### v4

# Causal Inference in Longitudinal Data under Unknown Interference

Abstract
In longitudinal studies where units are embedded in space or a social network, interference may arise, meaning that a unit’s outcome can depend on treatment histories of others. The presence of interference poses significant challenges for causal inference, particularly when the interference structure—how a unit’s outcome responds to others’ influences—is complex, heterogeneous, and unknown to researchers. This paper develops a general framework for identifying and estimating both direct and spillover effects of treatment histories under minimal assumptions about the interference structure. We define a class of policy-relevant causal estimands and show that they can be represented by a modified marginal structural model (MSM). Under the standard assumption of sequential exchangeability, these estimands are identifiable and can be estimated using inverse probability weighting (IPW). We derive conditions for consistency and asymptotic normality of the estimators and provide procedures for constructing Wald-type confidence intervals with valid coverage in large samples. The method’s utility is demonstrated through applications in both social science and biomedical settings.
Keywords: Causal Inference, Interference, Longitudinal Data, Marginal Structural Models, Sequential Exchangeability, Inverse Probability Weighting
1 Introduction
Empirical researchers are often interested in estimating the causal effects of treatment histories in longitudinal data. For example, public health researchers may seek to determine whether a history of smoking increases the likelihood of developing cardiovascular diseases, while social scientists want to understand how past experiences with government programs shape voting behavior. Under the standard assumption of sequential exchangeability—in either experimental or observational studies—these effects can be represented by a marginal structural model (MSM) and consistently estimated through inverse probability weighting (IPW) (Robins, Hernan and Brumback, 2000; Hernán and Robins, 2010).
However, the classic approach relies on the stable unit treatment value assumption (SUTVA), which rules out the possibility that the outcome of a unit is affected by treatment histories of others, a phenomenon known as spillover or interference (Cox, 1958). In contexts where units are embedded in geographic space or social networks, interference is likely present, raising questions on how researchers can describe and estimate the direct or spillover effects a treatment history generates. Most existing solutions assume a known interference structure, typically formalized as an “exposure mapping” that specifies how each unit’s outcome depends on the treatments received by a fixed set of others (Aronow and Samii, 2017; Ogburn et al., 2020). For example, in studies of second-hand smoke, it is often assumed that an individual’s health is influenced only by the average smoking behavior of their immediate contacts.
In many applications, nevertheless, the interference structure is complex, heterogeneous, and unknown to researchers. Individuals may be affected by second-hand smoke from socially or spatially distant others, depending on their unobserved daily activities. In such cases, which we refer to as “unknown interference,” exposure mappings may fail to capture the relevant sources of influence, undermining identification and estimation.
In this paper, we present a general framework to address this problem by integrating existing methods in longitudinal data analysis with recent advancements on causal inference under unknown interference (Sävje, Aronow and Hudgens, 2021; Wang et al., 2025). We introduce a novel estimand, the average marginalized response (AMR), which can be used to capture either the direct or spillover effects and remains well-defined even when the interference structure is unknown. The only requirement is the existence of a proximity metric between any two units (e.g., spatial distance, network path length, or shared group membership), which is necessary for indexing spillover effects but but imposes no restrictions on the form or complexity of the interference structure.
For a fixed proximity level $d\geq 0$, the AMR summarizes how, on average, the expected outcomes of a unit’s neighbors at that proximity level respond to the unit’s treatment history. As a function of treatment history, AMR can naturally be represented by a MSM, where linear combinations of the parameters reflect the causal effect of a unit’s treatment history on the outcomes of its neighbors at proximity level $d$. By evaluating AMRs and their associated MSMs across different values of $d$, researchers can examine how the effects of treatment histories vary over the proximity to one’s neighbors and distinguish the direct effect ($d=0$) from spillover effects ($d>0$) along both temporal and cross-unit dimensions.
We then demonstrate that both the AMR and the parameters in the associated MSM can be identified under the familiar assumption of sequential exchangeability. This identification result enables estimation of MSM parameters using IPW estimators, such as weighted least squares (WLS), where we regress the average outcome of a unit’s neighbors at the proximity level of $d$ on indicators of its own treatment history, weighted by the inverse probability of that history being observed. We prove that this estimator exhibits desirable large-sample properties—specifically, consistency and asymptotic normality—when either the influence of a unit’s treatment decays sufficiently quickly over proximity or neighborhood sizes grow slowly with the sample size, using recently developed central limit theorems for dependent data (Kojevnikov, Marmer and Song, 2021; Leung, 2022). In addition, we provide a heteroskedasticity and autocorrelation-consistent (HAC) variance estimator that is guaranteed to be conservative, enabling researchers to construct Wald-style confidence intervals with valid coverage in large samples.
We explore several extensions of the method, including evaluating the identification assumptions using placebo tests and allowing for diffusion or dependence in treatment assignment. We also discuss the implications of misspecifying the MSM and having measurement error in the proximity metric. We test the method’s performance via Monte Carlo simulations and demonstrate its application through two empirical studies. The first study examines whether wind turbine proposals generate spillover effects in space on political behavior among Canadian voters. The second investigates whether exposure to second-hand smoke through social connections can impact health outcomes in the Framingham Heart Study.
This paper contributes to several strands of the causal inference literature. The definition of the AMR extends the concept of marginalized causal effects (Hudgens and Halloran, 2008; Sävje, Aronow and Hudgens, 2021; Hu, Li and Wager, 2022; Wang et al., 2025) to longitudinal data. We clarify its connection to MSMs and show how it can capture not only direct effects (Papadogeorgou et al., 2020), but also a wide range of spillover effects. Compared to alternative approaches—particularly those based on exposure mappings (Ogburn et al., 2020; Jiang, Wallace and Thompson, 2023; Jetsupphasuk, Li and Hudgens, 2025)—our method requires only minimal knowledge of the interference structure and remains valid even when the effects from other units are complex and heterogeneous. It is also robust to dependence in confounders and contagion in outcomes and avoids the additional step to compute exposure probabilities (Aronow and Samii, 2017; Forastiere, Airoldi and Mealli, 2021).
The estimators we propose mirror the form of classical methods in longitudinal data analysis when interference is absent (Robins, 1986; Robins, Hernan and Brumback, 2000; Cole and Hernán, 2008) and rely on familiar identification assumptions. They are straightforward to implement for applied researchers across biomedical and social sciences and are accompanied by a complete procedure for statistical inference that we extend to estimates of spillover effects (Kojevnikov, Marmer and Song, 2021; Leung, 2022; Gao and Ding, 2023). By incorporating more information from treatment assignment, our approach complements methods built upon outcome models, which are commonly employed in both network and spatio-temporal analyses (Ogburn, Shpitser and Lee, 2020; Reich et al., 2021).
The rest of the paper is organized as follows: Section 2 describes the basic framework and defines the estimands. Section 3 discusses identification and introduces the estimators. Section 4 establishes large-sample properties of the proposed estimators and develops methods for conducting statistical inference. Section 5 explores possible extensions. Section 6 presents results from two empirical analyses. Section 7 concludes.
2 The Framework
2.1 Set Up
We focus on complete longitudinal data from either experiments or observational studies with $N$ units spanning over $T+1$ periods, where $N\gg T$ [^3] We discuss scenarios with a large $T$ in Section 5. Throughout the paper, uppercase letters denote random variables, lowercase letters represent their realizations, and boldface indicates vectors or matrices. Subscripts refer to specific units, while superscripts in parentheses indicate specific time periods. For each unit $i\in\mathcal{N}=\{1,2,\dots,N\}$ in period $t\in\{0,1,2,\dots,T\}$, we observe the treatment status $A_{i}^{(t)}\in\{0,1\}$ and a vector of time-varying confounders $\mathbf{L}_{i}^{(t)}$, which can also include time-invariant variables. For simplicity, we assume that $A_{i}^{(0)}=0$ for all $i$. The outcome $Y_{i}$ is measured in the final period $T$. We assume that units in the data are embedded in a fixed social network or geographic space, $\mathcal{G}_{N}$, with a time-invariant proximity metric $d_{ij}$ defined for every pair of units $i$ and $j$.
Figure 1: A DAG illustration
[…]
Notes: The DAG depicts two units, $\{i,j\}$, observed over two periods, $\{T-1,T\}$. Variables are represented by circles and causal paths by arrows. White circles denote treatment, gray circles represent time-varying confounders, and black circles correspond to outcomes. Red arrows illustrate the effect of a unit’s own treatment history, while blue arrows capture spillover effects due to interference. Black arrows indicate relationships influencing treatment assignment, and gray arrows represent other potential dependencies between variables under the assumptions outlined in Section 3 (e.g., causal mechanisms).
We adopt the potential outcomes framework and allow for an unknown interference structure. We represent unit $i$’s treatment history between period $1$ and period $T$ by $\mathbf{A}_{i}^{(1:T)}=\left(A_{i}^{(1)},A_{i}^{(2)},\dots,A_{i}^{(T)}\right)$. For any subset $\mathcal{S}\subseteq\mathcal{N}$, $\mathbf{A}_{\mathcal{S}}^{(1:T)}$ denotes the collection of $\mathbf{A}_{i}^{(1:T)}$ for $i\in\mathcal{S}$, an $(|\mathcal{S}|\times T)$-dimensional vector. Both the outcome variable $Y_{i}$ and time-varying confounders $\mathbf{L}_{i}^{(t)}$ are functions of the full treatment histories of all the $N$ units, $\mathbf{A}_{\mathcal{N}}^{(1:T)}$:
$Y_{i}=Y_{i}\left(\mathbf{A}_{\mathcal{N}}^{(1:T)}\right),\mathbf{L}_{i}^{(t)}=\mathbf{L}_{i}^{(t)}\left(\mathbf{A}_{\mathcal{N}}^{(1:T)}\right),$ (1)
where the functional forms are unknown to the researcher and may vary across units. Since $\mathbf{A}_{\mathcal{N}}^{(1:T)}$ is an $(N\times T)$-dimensional vector, there could be $2^{N\times T}$ different potential values for each $Y_{i}$ or $\mathbf{L}_{i}^{(t)}$, in contrast to $2^{T}$ values in the classic setting without interference. Figure 1 depicts the relationships across variables in our framework using a directed acyclic graph (DAG) with two units and two time periods. The arrows from $A_{i}^{(T-1)}$ and $A_{i}^{(T)}$ to $Y_{i}$ represent the causal effect of unit $i$’s own treatment history, while the arrows from $A_{j}^{(T-1)}$ and $A_{j}^{(T)}$ to $Y_{i}$ represent spillover effects due to interference.
An illustration
We provide a concrete example for concepts defined above using simulated data with $N=400$ and $T=2$. The units can be understood as either nodes in a social network (top-left plot of Figure 2) or tiles in a spatial raster (top-right plot of Figure 2). As shown in the bottom-left plot of Figure 2, there are four possible treatment histories: $\mathbf{A}_{i}^{(1:2)}\in\{(0,0),(0,1),(1,0),(1,1)\}$. Consider the outcome of an arbitrary unit $i$ (e.g., unit $42$). Without further restrictions, its value is jointly determined by the treatment histories of all the $400$ units: $Y_{42}=Y_{42}\left(\mathbf{A}_{\{1,2,\dots,400\}}^{(1:2)}\right)=Y_{42}\left(\mathbf{A}_{1}^{(1:2)},\mathbf{A}_{2}^{(1:2)},\dots,\mathbf{A}_{400}^{(1:2)}\right)$. [^4] Details of the data generating process are described in the Supplementary Material.
Figure 2: Simulated Data and Treatment Assignment
Notes: The top plots show the locations of the $400$ units from the simulation in either a social network (left) or a $20\times 20$ spatial raster object (right), where the colors indicate the outcome’s variation. Red colors on the left mark a node ($42$) and its second-degree neighbors in the network, while red circles on the right show the construction of $\Omega_{i}(d)$ around the spotted unit. The bottom-left plot depicts the treatment assignment history for the first $10$ units in the sample, while the bottom-right plot presents combinations of MSM parameters that capture various types of causal effects.
2.2 Estimands
Our goal is to understand how the treatment history of a unit influences its own outcome or the outcome of its neighbors in $\mathcal{G}_{N}$. Let $\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}$ represent the remaining part in $\mathbf{A}_{\mathcal{N}}^{(1:T)}$ after removing unit $j$’s treatment history between periods $1$ and $T$. Then, we use $Y_{i}\left(\mathbf{a}^{(1:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)=Y_{i}\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)$ to denote the potential outcome of unit $i$ when the treatment history of unit $j$ is fixed at $\mathbf{a}^{(1:T)}$ but all other units’ treatment histories are assigned at random. To construct meaningful estimands, we take the expectation of $Y_{i}\left(\mathbf{a}^{(1:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)$ over $\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}$ and other random variables that may affect $Y_{i}$, obtaining the marginalized potential outcome for unit $i$ when $j$’s treatment history is fixed:
$Y_{i;j}\left(\mathbf{a}^{(1:T)}\right)\coloneqq{\rm E}\,\left[Y_{i}\left(\mathbf{a}^{(1:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)\right].\vspace{-0.5em}$ (2)
This quantity can be defined for every pair of $i$ and $j$, including cases where $i=j$.
Next, for each unit $j$, we aggregate $Y_{i;j}\left(\mathbf{a}^{(1:T)}\right)$ over a pre-specified set of $j$’s neighbors denoted as $\Omega_{j}(d)$, where $d$ is a specific value of the proximity metric. This leads to a linear mapping $\mu\left(\cdot;\Omega_{j}(d)\right)$ from the non-random vector $\big{\{}Y_{i;j}\left(\mathbf{a}^{(1:T)}\right)\big{\}}_{i\in\mathcal{N}}$ to a real number:
$\mu\left(\big{\{}Y_{i;j}\left(\mathbf{a}^{(1:T)}\right)\big{\}}_{i\in\mathcal{N}};\Omega_{j}(d)\right)$ $\coloneqq\frac{\sum_{i=1}^{N}\mathbf{1}\{i\in\Omega_{j}(d)\}Y_{i;j}\left(\mathbf{a}^{(1:T)}\right)}{|\Omega_{j}(d)|},$ (3)
where $\mathbf{1}\{\cdot\}$ is the indicator function and $|\Omega_{j}(d)|=\sum_{i=1}^{N}\mathbf{1}\{i\in\Omega_{j}(d)\}$ represents the number of units in $\Omega_{j}(d)$. The form of $\Omega_{j}(d)$ can be chosen by the researcher based on the type of spillover effects under study. In social networks, a natural choice is the $d$th-degree neighbors of each $j$: $\Omega_{j}(d)\coloneqq\{i\in\mathcal{N}:d_{ij}=d\}$. In space, one option is the “circle average” introduced by Wang et al. (2025), where $d$ stands for any distance value and $\Omega_{i}(d)$ represents a “doughnut:” $\{i\in\mathcal{N}:d-\kappa<d_{ij}\leq d\}$, with $\kappa$ being a fixed constant. In either case, the quantity defined in (3) captures the expected response of unit $j$’s neighbors whose proximity to $i$ is (approximately) $d$ to $i$’s treatment history. With a slight abuse of terminology, we will refer to units belonging to $\Omega_{j}(d)$ as $j$’s $d$th-degree neighbors henceforth. When $d=0$, $\Omega_{j}(d)$ only includes unit $j$ itself, and the quantity boils down to $Y_{j;j}\left(\mathbf{a}^{(1:T)}\right)$, unit $j$’s expected outcome given its own treatment history.
In the top-left plot of Figure 2, all the second-degree neighbors ($d=2$) of unit $42$ are marked in red. In the top-right plot, a specific unit $j$ is highlighted in red, while red circles indicate the range of $\Omega_{j}(d)$ (in the form of doughnuts with $\kappa=1$) for $d\in\{0,1,2,3\}$. Both $d$ and $\Omega_{j}(d)$ can be adapted to specific contexts. For example, $d$ may take discrete values indicating whether two units are from the same street or district, or represent traffic accessibility or even cultural proximity between units. We discuss potential impacts of measurement error of $d$ in Section 5 and summarize choices of $d$ and $\Omega_{j}(d)$ in the Supplementary Material. When the context is clear, we represent the value of $\mu(\cdot;\Omega_{j}(d))$ with $\mu_{j}(\cdot;d)$.
It is worth noting that the definition of $\Omega_{j}(d)$ does not require any knowledge of the interference structure—specifically, the functional form of $Y_{i}\left(\mathbf{A}_{\mathcal{N}}^{(1:T)}\right)$. It remains well-defined as long as the proximity metric exists. Instead of explicitly studying how a unit’s outcome is influenced by the others—typically formalized through an exposure mapping—our framework examines the effects of a unit’s treatment history on a predefined set of neighbors. We do not assume that any $\Omega_{j}(d)$ or their collection captures all the influences generated by a unit. The form of $\Omega_{j}(d)$ is flexible and not unique. [^5] But it does impact statistical inference. We discuss this issue in Section 4. Overlap of $\Omega_{j}(d)$ across the units is permitted, hence each unit may contribute to multiple $\mu_{j}\left(\Big{\{}Y_{i;j}\left(\mathbf{a}^{(1:T)}\right)\Big{\}}_{i\in\mathcal{N}};d\right)$ values if it is influenced by more than one neighbor.
Finally, by taking the average of $\mu_{j}\left(\Big{\{}Y_{i;j}\left(\mathbf{a}^{(1:T)}\right)\Big{\}}_{i\in\mathcal{N}};d\right)$ across all units in the sample, we obtain the average marginalized response (AMR):
$\mu\left(\mathbf{a}^{(1:T)};d\right)\coloneqq\frac{1}{N}\sum_{j=1}^{N}\mu_{j}\left(\Big{\{}Y_{i;j}\left(\mathbf{a}^{(1:T)}\right)\Big{\}}_{i\in\mathcal{N}};d\right).\vspace{-0.5em}$ (4)
Given a pre-specified range $\mathcal{D}$, the AMR can be defined for each $d\in\mathcal{D}$ and enables researchers to assess how spillover effects vary with proximity without making assumptions about their magnitude. For instance, we could set $\mathcal{D}=\{0,1,2,3,4\}$ if the goal is to test whether spillover effects only extend to one’s third-degree neighbors. Moreover, the AMR at any $d$ depends only on the treatment history $\mathbf{a}^{(1:T)}$, as other factors affecting the outcome have been marginalized over. Thus, when $T$ is not too large, the relationship between any AMR and the corresponding treatment history can be represented by a saturated marginal structural model (MSM):
$\mu\left(\mathbf{a}^{(1:T)};d\right)=\mathbf{m}\left(\mathbf{a}^{(1:T)}\right)^{\prime}\boldsymbol{\beta}(d).\vspace{-1em}$ (5)
Here, $\mathbf{m}\left(\mathbf{a}^{(1:T)}\right)$ consists of indicators for the treatment status in each period and their interactions, and $\boldsymbol{\beta}(d)$ is a vector of parameters specific to $d$. When $d=0$, Equation (5) resembles a conventional MSM, where the parameters describe the direct effects of a unit’s treatment histories on its own expected outcome. When $d>0$, $\boldsymbol{\beta}(d)$ captures spillover effects those treatment histories generate on the expected outcome of the unit’s $d$th-degree neighbors. [^6] As in a conventional MSM, these quantities include the effects mediated through the time-varying confounders, represented by gray arrows in Figure 1. $\{\boldsymbol{\beta}(d)\}_{d\in\mathcal{D}}$ thus provides a comprehensive view of both direct and spillover effects in the sample.
Similar to other marginalized estimands in the literature (Hudgens and Halloran, 2008; Sävje, Aronow and Hudgens, 2021; Hu, Li and Wager, 2022; Wang et al., 2025), these causal quantities reflect the effect of manipulating a single unit’s treatment history under the current assignment policy, rather than contrasting hypothetical scenarios where all units receive specific treatments. They are inherently sample-specific (though we omit $\mathcal{N}$ in the subscript for simplicity) and “descriptive” of the status quo, not “prescriptive” for identifying the optimal policy (Kennedy, 2019). As such, $\boldsymbol{\beta}(d)$ does not represent structural parameters invariant to changes in the assignment policy. Nevertheless, it remains meaningful under the prevailing policy, capturing all causal pathways through which treatment influences outcomes, including direct effects, spillover effects, changes in time-varying confounders, and shifts in treatment distributions over time.
Table 1: Examples of AMR
$\mathbf{a}^{1:2}$ $(0,0)$ $(0,1)$ $(1,0)$ $(1,1)$
$AMR$ $\mu\left((0,0);d\right)$ $\mu\left((0,1);d\right)$ $\mu\left((1,0);d\right)$ $\mu\left((1,1);d\right)$
Associated $\boldsymbol{\beta}(d)$ $\beta_{0}(d)$ $\beta_{0}(d)+\beta_{2}(d)$ $\beta_{0}(d)+\beta_{1}(d)$ $\beta_{0}(d)+\beta_{1}(d)+\beta_{2}(d)+\beta_{3}(d)$
An illustration (continued)
In the example introduced previously, we can define a series of AMRs for each value of $\mathbf{a}^{(1:T)}$ and $d\in\mathcal{D}=\{0,1,\dots,7\}$, where $d$ is measured by the shortest length path in the social network and the Euclidean distance in the spatial raster. $\Omega_{j}(d)$ is defined as $j$’s $d$th-degree neighbors in the former case and a doughnut surrounding $j$ with $\kappa=1$ in the latter, as depicted in the top plots of Figure 2. The top row of Table 2 presents AMRs defined under each of the four possible treatment histories, which can be summarized by the following MSM at any $d$:
$\mu\left(\mathbf{a}^{(1:T)};d\right)=\beta_{0}(d)+\beta_{1}(d)a^{(1)}+\beta_{2}(d)a^{(2)}+\beta_{3}(d)a^{(1)}a^{(2)},\vspace{-1em}$
with $\mathbf{m}\left(\mathbf{a}^{(1:T)}\right)=\left(1,a^{(1)},a^{(2)},a^{(1)}a^{(2)}\right)^{\prime}$. Since this model is saturated, the combination of its parameters can recover each of the AMRs, as presented in the bottom row of Table 2. These parameters can also be used to construct the causal contrast between two treatment histories. For instance, the effect of $(1,1)$ relative to $(0,0)$ equals $\beta_{1}(d)+\beta_{2}(d)+\beta_{3}(d)$ in the model.
We specify an effect function that emanates from each unit and declines in $d$ exponentially. Effects received by each unit are additive across other units in the sample and amplified by an idiosyncratic constant that represents treatment effect heterogeneity (depicted by varying colors in the plot). The effects in each period carry over into the next, declining by $50\%$ in magnitude. In period $t$, the probability for unit $i$ to receive the treatment depends on its treatment status and the value of a time-varying confounder in the previous period. We repeat the simulation $1,000$ times and approximate each $Y_{i;j}\left(\mathbf{a}^{(1:T)}\right)$ by the average of $Y_{i}\left(\mathbf{a}^{(1:T)};\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)$ over the simulations. The AMRs are then constructed in line with their definition in Equation (4). We plot several linear combinations of the MSM parameters against proximity levels in the bottom-right plot of Figure 2. As shown, each curve captures how the effects generated by a specific treatment history, relative to the benchmark $(0,0)$, vary with proximity.
3 Identification and Estimation
3.1 Identification Assumptions
To identify the AMR and parameters in the associated MSM, we impose the assumption of sequential exchangeability:
Assumption 1 (Sequential exchangeability)
$\Big{\{}Y_{i}\left(\mathbf{a}^{(1:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right),\mathbf{L}_{i}^{(t+1)}\left(\mathbf{a}^{(1:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)\Big{\}}\perp A_{j}^{(t)}\mid\left(\mathbf{A}_{j}^{(1:(t-1))},\mathbf{L}_{j}^{(1:t)}\right).\vspace{-1em}$
for any $i$, $j$, $t$, and $\mathbf{a}^{(1:T)}$.
Assumption 1 states that for any unit $j$, its treatment status in period $t$, $A_{j}^{(t)}$, is independent to the distribution of any unit’s potential outcomes or the potential values of time-varying confounders in period $t+1$—rather than just its own—conditional on its history of treatment assignment and time-varying confounders. The assumption reduces to the familiar version in the absence of interference, where $Y_{i}\left(\mathbf{a}^{(1:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)=Y_{i}\left(\mathbf{a}^{(1:T)}\right)$, $\mathbf{L}_{i}^{(t+1)}\left(\mathbf{a}^{(1:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)=\mathbf{L}_{i}^{(t+1)}\left(\mathbf{a}^{(1:T)}\right)$, and units are independent from each other. Assumption 1 is violated if the treatment assignment of unit $j$ depends on the history of another unit, thereby excluding any form of diffusion or dependence in treatment across units (VanderWeele and An, 2013; Sävje, Aronow and Hudgens, 2021). This requirement is naturally satisfied in experiments and observational studies where a unit’s treatment status is determined independently, without reference to others, and commonly adopted in the literature (e.g., Liu, Hudgens and Becker-Dreps, 2016). In Section 5, we explore ways to extend the method to settings where treatment diffusion exists or treatment assignment is correlated across units within the same period.
Assumption 1 tends to be weaker than the assumptions required for exposure mapping methods, which demand that the level of exposure for any unit $i$—typically a function of treatment histories within a subset of $i$’s neighbors—is independent of $i$’s potential outcomes conditional on its history (Forastiere, Airoldi and Mealli, 2021). This requirement fails in cases where $Y_{i}$ is influenced by $\mathbf{L}_{j}^{(T-1)}$, a phenomenon known as “contextual factors” when $\mathbf{L}_{j}^{(T-1)}$ contains time-invariant variables (VanderWeele and An, 2013; Egami and Tchetgen Tchetgen, 2024) or “contagion” when it incorporates past values of the outcome (Ogburn and VanderWeele, 2017). In contrast, Assumption 1 remains valid even in these scenarios.
We assume that positivity always holds. Denoting $\left(\mathbf{A}_{j}^{(1:(t-1))},\mathbf{L}_{j}^{(1:t)}\right)$ as $\mathbf{V}_{j}^{(t)}$, the assumption can be stated as:
Assumption 2 (Positivity)
For any $j$ and $t$, there exists some $\eta>0$ such that
$1-\eta<P\left(A_{j}^{(t)}=1\mid\mathbf{V}_{j}^{(t)}\right)<\eta.\vspace{-0.5em}$
With Assumptions 1 and 2 satisfied, the propensity score for unit $j$’s treatment history $\mathbf{a}^{(1:T)}$ can be expressed as $e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)=\prod_{s=1}^{T}P\left(A_{j}^{(s)}=a^{(s)}\mid\mathbf{V}_{j}^{(s)}\right)$. These probabilities are known in experimental settings and need to be estimated from data in observational studies.
Next, we impose the restriction that any time-varying confounder is influenced only by treatment histories up to the present and not by future treatment statuses. This assumption, commonly referred to as no anticipation (e.g., Hernán and Robins, 2010), is formally stated as follows:
Assumption 3 (No anticipation)
For any treatment history $\mathbf{a}^{(1:T)}$ and $\tilde{\mathbf{a}}^{(1:T)}$ and any period $t$,
$\mathbf{L}_{i}^{(t+1)}\left(\mathbf{a}^{(1:T)}\right)=\mathbf{L}_{i}^{(t+1)}\left(\tilde{\mathbf{a}}^{(1:T)}\right)\text{ if }\mathbf{a}^{(1:t)}=\tilde{\mathbf{a}}^{(1:t)}.\vspace{-1em}$
Although Assumption 3 is not essential for causal identification, it simplifies estimation and facilitates placebo tests (see Section 5 for details). In Section 5, we discuss possible relaxations of this assumption. We present structural equations that describe the relationships among variables under Assumptions 13 and discuss their connection to the framework in Ogburn et al. (2020) in the Supplementary Material.
3.2 Estimator
The key insight of the paper is that assumptions introduced in the previous subsection are sufficient for identifying the AMR at any $d$ and the parameters in the associated MSM. Define the transformed outcome of unit $j$ at $d$ as
$\mu_{j}(d)=\mu\left(\big{\{}Y_{i}\big{\}}_{i\in\mathcal{N}};\Omega_{j}(d)\right)=\frac{\sum_{i=1}^{N}\mathbf{1}\{i\in\Omega_{j}(d)\}Y_{i}}{|\Omega_{j}(d)|},\vspace{-1em}$ (6)
which represents the output of the same mapping $\mu(\cdot;\Omega_{j}(d))$, but applied to the vector of observed outcomes rather than marginalized potential outcomes. We then establish the following identification result:
Proposition 1
Under Assumptions 1-3,
$\mu\left(\mathbf{a}^{(1:T)};d\right)=\frac{1}{N}\sum_{j=1}^{N}{\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\}\mu_{j}(d)}{e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)}\right].\vspace{-1em}$ (7)
Since each parameter in the MSM is a linear combination of the AMRs, they can be similarly identified from the data. The proofs are provided in the Supplementary Material.
As in the classic case, these identification results suggest that $\boldsymbol{\beta}(d)$ can be estimated using IPW. Specifically, we rely on the weighted least squares (WLS) algorithm to obtain:
$\hat{\boldsymbol{\beta}}(d)={\rm arg}\,\min_{\boldsymbol{\beta}(d)}\sum_{j=1}^{N}w_{j}\left(\mu_{j}(d)-\mathbf{m}\left(\mathbf{A}_{j}^{(1:T)}\right)^{\prime}\boldsymbol{\beta}(d)\right)^{2},\vspace{-1em}$ (8)
where $w_{j}=\frac{\prod_{s=1}^{T}P\left(A_{j}^{(s)}\mid\mathbf{A}_{j}^{(1:(s-1))}\right)}{e\left(\mathbf{A}_{j}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)}$ is the stabilized weight (Cole and Hernán, 2008). If $e\left(\mathbf{A}_{j}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)$ is unknown, as in an observational study, then we replace it with an estimate in $w_{j}$. In the following section, we show that this estimator possesses desirable large-sample properties, such as consistency and asymptotic normality, under mild restrictions on the dependence across units caused by interference.
Beyond its validity under unknown interference, the proposed method offers two other advantages over approaches based on exposure mapping. First, it only requires researchers to estimate the probability of receiving treatment, rather than the probability of attaining a specific level of exposure, which often involves computing a convolution. [^7] For instance, in a common exposure mapping where exposure is defined as the average of treatment statuses among a unit’s first-degree neighbors, computing this probability is equivalent to deriving the distribution of the sum of multiple random variables. Moreover, the exposure probabilities can be close to 0 or 1 even when Assumption 2 is satisfied. Second, because identification relies on Assumption 1, $e\left(\mathbf{A}_{j}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)$ does not need to account for variables influencing the treatment statuses of a unit’s neighbors, such as contextual factors or sources of “spatial confounding” (Papadogeorgou, Choirat and Zigler, 2019; Reich et al., 2021). Steps 1-4 in Algorithm 1 summarize the estimation procedure for the proposed method. We provide justifications for Steps 4 and 5 in the following section.
1 for $d\in\mathcal{D}$ do
2       for $j\in\mathcal{N}$ do
3             Construct the transformed outcome $\mu_{j}(d)$ based on Equation (6).
4            Compute the probability for $j$’s treatment history to occur and the corresponding weight $w_{j}=\frac{\prod_{s=1}^{T}P\left(A_{j}^{(s)}\mid A_{j}^{(1:(s-1))}\right)}{e\left(\mathbf{A}_{j}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)}$, estimating $e\left(\mathbf{A}_{j}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)$ if unknown.
5
6      Estimate $\boldsymbol{\beta}(d)$ by the WLS algorithm presented in Equation (8).
7      Estimate $\hat{\boldsymbol{\beta}}(d)$’s variance, $\Sigma_{N}(d)$, using the heteroskedesticity and auto-correlation consistent (HAC) variance estimator in Equation (10).
8      Construct Wald-type confidence intervals using critical values from the standard normal distribution.
9      (Optional) Test identification assumptions by applying the estimator to a placebo outcome.
Algorithm 1 Implementation of the Proposed Method
4 Large-Sample Theory
To derive the large-sample properties of our estimator, we consider a sequence $\{\mathcal{G}_{N}\}_{N>0}$ in which $\mathcal{N}_{N}$ units are embedded, with $N\to\infty$. Define the vector of transformed outcomes at proximity level $d$: $\boldsymbol{\mu}(d)=(\mu_{1}(d),\mu_{2}(d),\dots,\mu_{N}(d))^{\prime}$, the vector of outcomes: $\mathbf{Y}=(Y_{1},Y_{2},\dots,Y_{N})^{\prime}$, and the $N\times N$ matrix: $\mathbf{D}(d)=\Big{\{}\frac{\mathbf{1}\{i\in\Omega_{j}(d)\}}{|\Omega_{j}(d)|}\Big{\}}_{N\times N}$.
From Equation (6), we can see that $\boldsymbol{\mu}(d)=\mathbf{D}(d)\mathbf{Y}$. Next, define the matrix of regressors for the MSM as: $\mathbf{M}=\left(\mathbf{m}\left(\mathbf{A}_{1}^{(1:T)}\right),\mathbf{m}\left(\mathbf{A}_{2}^{(1:T)}\right),\dots,\mathbf{m}\left(\mathbf{A}_{N}^{(1:T)}\right)\right)^{\prime}$, and the diagonal weight matrix: $\mathbf{W}=diag\left(w_{1},w_{2},\dots,w_{N}\right)$. We know that estimates from the WLS algorithm can be expressed as
$\hat{\boldsymbol{\beta}}(d)=$ $\left(\mathbf{M}^{\prime}\mathbf{W}\mathbf{M}\right)^{-1}\left(\mathbf{M}^{\prime}\mathbf{W}\boldsymbol{\mu}(d)\right)=\left(\mathbf{M}^{\prime}\mathbf{W}\mathbf{M}\right)^{-1}\left(\mathbf{M}^{\prime}\mathbf{W}\mathbf{D}(d)\mathbf{Y}\right)$
$=$ $\left(\mathbf{M}^{\prime}\mathbf{W}\mathbf{M}\right)^{-1}\left(\mathbf{Z}^{\prime}(d)\mathbf{Y}\right)=\left(\frac{1}{N}\sum_{i=1}^{N}w_{i}\mathbf{m}\left(\mathbf{A}_{i}^{(1:T)}\right)\mathbf{m}^{\prime}\left(\mathbf{A}_{i}^{(1:T)}\right)\right)^{-1}\left(\frac{1}{N}\sum_{i=1}^{N}\mathbf{Z}_{i}(d)Y_{i}\right),\vspace{-1em}$
where $\mathbf{Z}^{\prime}(d)=\left(\mathbf{Z}_{1}(d),\mathbf{Z}_{2}(d),\dots,\mathbf{Z}_{N}(d)\right)=\mathbf{M}^{\prime}\mathbf{W}\mathbf{D}(d)$. Thus, the estimates take the form of linear combinations of the outcome $Y_{i}$.
Since $\mathbf{A}_{i}^{(1:T)}$ is conditionally independent across units, we can show the convergence of the first term, $\frac{1}{N}\sum_{i=1}^{N}w_{i}\mathbf{m}\left(\mathbf{A}_{i}^{(1:T)}\right)\mathbf{m}^{\prime}\left(\mathbf{A}_{i}^{(1:T)}\right)$, under standard regularity conditions. For the second term, we know that $\mathbf{Z}_{i}(d)=\sum_{j=1}^{N}\frac{\mathbf{1}\{j\in\Omega_{i}(d)\}\mathbf{m}\left(\mathbf{A}_{j}^{(1:T)}\right)w_{j}}{|\Omega_{j}(d)|}$, which is a weighted average of the MSM regressors across $i$’s neighbors in $\Omega_{i}(d)$. $\mathbf{Z}_{i}(d)$ and $\mathbf{Z}_{j}(d)$ are independent if $\Omega_{i}(d)$ and $\Omega_{j}(d)$ do not overlap. [^8] Suppose $\Omega_{j}(d)\coloneqq\{i\in\mathcal{N}:d_{ij}=d\}$, then this is true when $d_{ij}>2d$. Therefore, the large-sample performance of $\hat{\boldsymbol{\beta}}(d)$ is primarily driven by the dependence in $Y_{i}$, which may be caused by either interference or correlated contextual factors.
4.1 Asymptotic Distribution
We restrict the dependence in $Y_{i}$ by assuming approximate neighborhood interference (ANI) introduced in Leung (2022). Let unit $i$’s $p$-neighborhood in $\mathcal{G}_{N}$ be defined as $\mathcal{N}_{\mathcal{G}_{N}}(i,p)$, the set of units whose proximity to $i$ does not exceed $p$. We denote an independent copy of the treatment history outside $i$’s $p$-neighborhood as $\mathbf{A}^{(1:T)}_{i,p}=\left(\mathbf{A}^{(1:T)}_{\mathcal{N}_{\mathcal{G}_{N}}(i,p)},\tilde{\mathbf{A}}^{(1:T)}_{\mathcal{N}_{N}\setminus\mathcal{N}_{\mathcal{G}_{N}}(i,p)}\right)$. We then impose the following assumption:
Assumption 4 (Approximate Neighborhood Interference)
$\max_{i\in\mathcal{N}_{N}}E\Bigg{|}Y_{i}\left(\mathbf{A}^{(1:T)}\right)-Y_{i}\left(\mathbf{A}^{(1:T)}_{i,p}\right)\Bigg{|}\leq\theta_{N,p},$
where $\sup_{N}\theta_{N,p}\rightarrow 0$ as $p\rightarrow\infty$.
This assumption implies that for any bounded Lipschitz function $L(\cdot)$, the covariance between $L\left(\mathbf{Z}_{i}(d)Y_{i}\right)$ and $L\left(\mathbf{Z}_{j}(d)Y_{j}\right)$ can be bounded by a transformation of $\theta_{N,p}$. This property, known as $\psi$-dependence (Doukhan and Louhichi, 1999; Kojevnikov, Marmer and Song, 2021), ensures that the dependence between outcomes decays as the proximity between units increases, even though it may never disappear.
To describe the expansion rate of $\mathcal{N}_{\mathcal{G}_{N}}(i,p)$ as $N$ grows, we introduce the following two quantities.
$M_{N}(q,k)=\frac{1}{N}\sum_{i=1}^{N}\left|\mathcal{N}_{\mathcal{G}_{N}}(i,q)\right|^{k},\vspace{-1em}$
and
$\mathcal{H}_{N}(p,q)=\{(i,j,i^{\prime},j^{\prime}):j^{\prime}\in\mathcal{N}_{\mathcal{G}_{N}}(i,q),\,j^{\prime}\in\mathcal{N}_{\mathcal{G}_{N}}(j,q),\,d_{\{i,i^{\prime}\},\{j,j^{\prime}\}}=p\}.$
These metrics help characterize how the structure of the network or spatial proximity affects the dependence between units as $N$ increases. In particular, $M_{N}(q,k)$ describes higher-order moments of neighborhood sizes, while the set $\mathcal{H}_{N}(p,q)$ contains neighborhoods of size $q$ that are exactly $p$ away from one another.
For simplicity, we assume that the potential outcomes of $Y_{i}$ are always bounded:
Assumption 5 (Bounded potential outcomes)
There exists a constant $\tilde{y}<\infty$ such that
$|Y_{i}\left(\mathbf{a}^{(1:T)}\right)|\leq\tilde{y}\vspace{-1em}$
for all units $i$ and any treatment history $\mathbf{a}^{(1:T)}$.
We discuss potential relaxations of this assumption in Section 5. Additionally, we assume that the propensity score can be consistently estimated at a parametric rate (root-$N$ or faster) in observational studies:
Assumption 6 (Consistent estimation of the propensity score)
Either the propensity score is known, or there exists an estimator $\hat{e}\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{A}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)$ such that
$\sqrt{N}\left(\hat{e}\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{A}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)-e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{A}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)\right)=O_{P}\left(1\right).$
[…]
Theorem 1
Under Assumptions 1-6, if there exist $\epsilon>0$ and a sequence $\{q_{N}\}$ that grows with $N$ such that
$\frac{N^{\frac{3}{2}}}{\Sigma_{N}^{\frac{1}{2}}(d)}\tilde{\theta}_{N,q_{N}}^{1-\epsilon}(d)\to 0,\frac{1}{\sqrt{N}\Sigma_{N}^{\frac{3}{2}}(d)}M_{N}(q_{N},2)\to 0,\frac{1}{N^{2}\Sigma_{N}^{2}(d)}\sum_{p=0}^{N}|\mathcal{H}_{N}(p,q_{N})|\tilde{\theta}_{N,p}^{1-\epsilon}(d)\to 0,\vspace{-1em}$
then $\left(\Sigma_{N}(d)\right)^{-\frac{1}{2}}\sqrt{N}\left(\hat{\boldsymbol{\beta}}(d)-\boldsymbol{\beta}(d)\right)\to N\left(0,\mathbf{I}\right)$ as $N\to\infty$.
Intuitively, these additional conditions impose constraints on the expansion rate of $\mathcal{G}_{N}$, the decay in the depencence between two units, and their interactions. Consistency and asymptotic normality can be achieved when either neighborhoods in $\mathcal{G}_{N}$ grow slowly or the dependence between a unit and its remote neighbors declines sufficiently fast along the sequence $\{q_{N}\}_{N>0}$. In a geographic space, it is often reasonable to assume that two units $i$ and $j$ are no longer interfering with each other if their proximity is larger than a threshold $\bar{p}$. In this case, $\theta_{N,p}(d)=0$ if $p>\bar{p}$, and conditions 1 and 3 hold if $\Sigma_{N}(d)=O_{P}(N)$. Additionally, $\left|\mathcal{N}_{\mathcal{G}_{N}}(i,q_{N})\right|=O\left(q_{N}^{2}\right)$, hence $M_{N}(q_{N},2)=O\left(q_{N}^{4}\right)$. Condition 2 is satisfied if $q_{N}=o\left(N^{\frac{1}{2}}\right)$, consistent with results in Ogburn et al. (2020) and Wang et al. (2025). In a network where neighborhoods expand at a polynomial rate, $\left|\mathcal{N}_{\mathcal{G}_{N}}(i,q_{N})\right|=O\left(q_{N}^{k}\right)$, these conditions are satisfied if $\Sigma_{N}(d)=O_{P}(N)$, $q_{N}=o\left(N^{\frac{1}{k}}\right)$, and $\theta_{N,p}(d)=\frac{1}{p^{k+2}}$.
4.2 Variance Estimation
Existing research has relied on HAC variance estimators for statistical inference under interference in both social network (Leung, 2022) and spatial settings (Conley, 1999; Wang et al., 2025). Such estimators require researchers to specify a bandwidth $\bar{d}_{N}$ that grows with $N$ and account for the covariance between the outcomes of units $i$ and $j$ only if $d_{ij}\leq\bar{d}_{N}$. We can re-weigh neighbors of unit $i$ within the bandwidth using a kernel function. In theory, $\bar{d}_{N}$ should be chosen to match the sequence $\{q_{N}\}_{N>0}$. In practice, it is often determined by some rules of thumb. For instance, we can set set $\bar{d}_{N}=\lfloor\max\{2\bar{p},2d\}\rfloor$ when interference disappears beyond the proximity level $\bar{p}$. In social networks, Leung (2022) suggests the following criterion:
$\bar{d}_{N}=\lfloor\max\{d_{N}^{*},2d\}\rfloor,\text{ with }d_{N}^{*}=\begin{cases}\frac{1}{2}\mathcal{L}\left(\mathcal{G}_{N}\right)\text{ if }\mathcal{L}\left(\mathcal{G}_{N}\right)<2\frac{\log N}{\log\delta\left(\mathcal{G}_{N}\right)}\\ \mathcal{L}^{\frac{1}{3}}\left(\mathcal{G}_{N}\right)\text{ otherwise}\end{cases},\vspace{-1em}$
where $\delta\left(\mathcal{G}_{N}\right)$ and $\mathcal{L}\left(\mathcal{G}_{N}\right)$ denote the average degree and average path length in $\mathcal{G}_{N}$, respectively. Researchers should examine the robustness of the variance estimates by using various $\bar{d}_{N}$.
Directly applying these HAC variance estimators to our setting presents two challenges. First, the bandwidth is typically chosen based on the dependence in the original outcome $Y_{i}$, whereas our WLS estimator is constructed using the transformed outcome $\mu_{j}(d)$. Determining an appropriate bandwidth for $\mu_{j}(d)$ from $\bar{d}_{N}$ is not straightforward. Fortunately, as illustrated at the beginning of this section, the estimator $\hat{\boldsymbol{\beta}}(d)$ can be expressed as a linear combination of $Y_{i}$, which allows us to perform inference at the level of the original outcome. Denoting the residual from the WLS estimator for unit $i$ as $\hat{\varepsilon}_{i}(d)$, we can show that
$\sum_{i=1}^{N}w_{i}\mathbf{m}\left(\mathbf{A}_{i}^{(1:T)}\right)\hat{\varepsilon}_{i}(d)=\sum_{i=1}^{N}\hat{\boldsymbol{\varepsilon}}_{Yi}(d),$
[…]
Let $\hat{\boldsymbol{\varepsilon}}_{Y}(d)=(\hat{\boldsymbol{\varepsilon}}_{Y1}(d),\hat{\boldsymbol{\varepsilon}}_{Y2}(d),\dots,\hat{\boldsymbol{\varepsilon}}_{YN}(d))^{\prime}$, and consider the following matrix of weights based on the uniform kernel: [^10] Both Leung (2022) and Gao and Ding (2023) find that the uniform kernel performs better than alternatives in practice. $\mathbf{K}_{\bar{d}_{N}}=\big{\{}\mathbf{1}\{d_{ij}\leq\bar{d}_{N}\}\big{\}}_{N\times N}$. Then, the HAC variance estimator in our setting takes the form:
$\hat{\Sigma}(d)=N\left(\mathbf{M}^{\prime}\mathbf{W}\mathbf{M}\right)^{-1}\left(\hat{\boldsymbol{\varepsilon}}^{\prime}_{Y}(d)\mathbf{K}_{\bar{d}_{N}}\hat{\boldsymbol{\varepsilon}}_{Y}(d)\right)\left(\mathbf{M}^{\prime}\mathbf{W}\mathbf{M}\right)^{-1}.$ (9)
The second challenge is that the variance estimator may not even be conservative for $\Sigma_{N}(d)$ due to the covariance in treatment effects when interference is present, as demonstrated in both Leung (2022) and Wang et al. (2025). We solve this problem by adopting the proposal in Gao and Ding (2023). Since $\mathbf{K}_{\bar{d}_{N}}$ is symmetric, it admits an eigenvalue decomposition of the form $\mathbf{Q}_{N}\Lambda_{N}\mathbf{Q}_{N}$, where $\mathbf{Q}_{N}$ is the matrix of eigenvectors and $\Lambda_{N}$ is a diagonal matrix of eigenvalues. Gao and Ding (2023) propose replacing $\mathbf{K}_{\bar{d}_{N}}$ in Equation (9) with $\mathbf{K}_{\bar{d}_{N}}^{+}=\mathbf{Q}_{N}\max\{\Lambda_{N},0\}\mathbf{Q}_{N}$, with the maximum taken element-wise. The modified HAC variance estimator can be expressed as
$\hat{\Sigma}^{+}(d)=N\left(\mathbf{M}^{\prime}\mathbf{W}\mathbf{M}\right)^{-1}\left(\hat{\boldsymbol{\varepsilon}}^{\prime}_{Y}(d)\mathbf{K}_{\bar{d}_{N}}^{+}\hat{\boldsymbol{\varepsilon}}_{Y}(d)\right)\left(\mathbf{M}^{\prime}\mathbf{W}\mathbf{M}\right)^{-1}.$ (10)
To establish the large-sample properties of the variance estimator, we define the following two quantities:
$\mathcal{N}_{\mathcal{G}_{N}}^{\partial}(i,p)=\{j\in\mathcal{N}_{N}:d_{ij}=p\},\mathcal{J}_{N}\left(p,\bar{d}_{N}\right)=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathbf{1}\{d_{ij}=p\}\sum_{k=1}^{N}\left|K_{\bar{d}_{N},ik}\right|\sum_{l=1}^{N}\left|K_{\bar{d}_{N},jl}\right|.\vspace{-1em}$
Here, $\mathcal{N}_{\mathcal{G}_{N}}^{\partial}(i,p)$ represents the set of unit $i$’s neighbors whose proximity to $i$ is exactly $p$. This set coincides with $\Omega_{i}(p)$ when the latter is defined as unit $i$’s $p$th-degree neighbors. $\mathcal{J}_{N}\left(p,\bar{d}_{N}\right)$ can also be expressed as $\{(i,j,i^{\prime},j^{\prime}):j^{\prime}\in\mathcal{N}_{\mathcal{G}_{N}}(i,q),\,j^{\prime}\in\mathcal{N}_{\mathcal{G}_{N}}(j,q),\,d_{ij}=p\}$, which generalizes $\mathcal{H}_{N}(p,q)$ defined above.
We further define $\mathbf{K}_{\bar{d}_{N}}^{-}=\mathbf{Q}_{N}\min|\{\Lambda_{N},0\}|\mathbf{Q}_{N}$. It follows that $\mathbf{K}_{\bar{d}_{N}}^{+}=\mathbf{K}_{\bar{d}_{N}}+\mathbf{K}_{\bar{d}_{N}}^{-}$ and $\quad M_{N}\left(\bar{d}_{N},k\right)=\frac{1}{N}\sum_{i=1}^{N}\left|\sum_{j=1}^{N}K_{\bar{d}_{N},ij}\right|^{k}$, where $K_{\bar{d}_{N},ij}$ denotes the $(i,j)$th entry of $\mathbf{K}_{\bar{d}_{N}}$. Analogously, we define $M_{N}^{-}\left(\bar{d}_{N},k\right)$ and $\mathcal{J}_{N}^{-}\left(p,\bar{d}_{N}\right)$ as counterparts to $M_{N}\left(\bar{d}_{N},k\right)$ and $\mathcal{H}_{N}\left(p,\bar{d}_{N}\right)$, respectively, using the entries of $\mathbf{K}_{\bar{d}_{N}}^{-}$. We then have the following result:
Theorem 2
Under Assumptions 1-6, if
$\sum_{p=0}^{N}\mathcal{N}_{\mathcal{G}_{N}}^{\partial}(i,p)\tilde{\theta}_{N,\bar{d}_{N}}^{1-\varepsilon}(d)=O(1),M_{N}\left(\bar{d}_{N},1\right)=O(N^{\frac{1}{2}}),M_{N}^{-}\left(\bar{d}_{N},1\right)=O(N^{\frac{1}{2}}),$
$M_{N}\left(\bar{d}_{N},2\right)=O(N),M_{N}^{-}\left(\bar{d}_{N},2\right)=O(N),\sum_{p=0}^{N}\mathcal{J}_{N}(p,\bar{d}_{N})\tilde{\theta}_{N,\bar{d}_{N}}(d)=O(N^{2}),$
and $\text{and }\sum_{p=0}^{N}\mathcal{J}^{-}_{N}(p,\bar{d}_{N})\tilde{\theta}_{N,\bar{d}_{N}}(d)=O(N^{2}),\vspace{-1em}$
then $\lim_{N\to\infty}\hat{\Sigma}^{+}(d)\succeq\Sigma(d)$.
5 Extensions
We explore several extensions of the proposed method in this section, with technical details provided in the Supplementary Material.
Placebo Tests
It is worth noting that causal identification only requires Assumption 1 to hold for potential outcomes rather than potential values of $\mathbf{L}_{i}^{(t)}$. However, the broader formulation, combined with Assumption 3, allows researchers to evaluate the assumption’s validity through placebo tests when certain variables in $\mathbf{L}_{i}^{(t)}$ are not expected to be correlated with treatment. For example, consider units with either of the two treatment histories, $(0,0)$ and $(0,1)$, in our simulated data, and suppose that $\mathbf{V}_{j}^{(t)}$ only includes variables from period $t-1$. Suppose that $\mathbf{L}_{i}^{(1)}\perp A_{j}^{(2)}\mid\mathbf{V}_{j}^{(2)}=\left(A_{j}^{(1)},\mathbf{L}_{j}^{(2)}\right)$. Then, we should expect to find a null effect of $A_{j}^{(2)}$ on the transformed outcome constructed from $\mathbf{L}_{i}^{(1)}$ using the WLS estimator. We refer to such variables in $\mathbf{L}_{i}^{(t)}$ as placebo outcomes in Algorithm 1.
Relaxed Assumptions
We consider two relaxations of Assumption 1. The first allows a unit’s treatment status in period $t$ to be influenced by the histories of nearby units, that is, treatment diffusion is permitted. We formalize the modified assumption as follows:
Assumption 7 (Sequential exchangeability with treatment diffusion)
$\Big{\{}Y_{i}\left(\mathbf{a}^{(1:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right),\mathbf{L}_{i}^{(,t+1)}\left(\mathbf{a}^{(1:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)\Big{\}}\perp A_{j}^{(t)}\mid\left(\mathbf{A}_{\{j\}\cup\mathcal{N}_{\mathcal{G}_{N}}(j,p)}^{(1:(t-1))},\mathbf{L}_{\{j\}\cup\mathcal{N}_{\mathcal{G}_{N}}(j,p)}^{(1:(t-1))}\right).\vspace{-1em}$
for any $i$, $j$, $t$, $\mathbf{a}^{(1:T)}$, and known $p$.
Here, $\mathcal{N}_{\mathcal{G}_{N}}(j,p)$ denotes the set of units whose proximity to $j$ does not exceed $p$, as defined in Section 4.1. Under Assumption 7, the variables $\mathbf{A}_{\mathcal{N}_{\mathcal{G}_{N}}(j,p)}^{(1:(t-1))}$ and $\mathbf{L}_{\mathcal{N}_{\mathcal{G}_{N}}(j,p)}^{(1:(t-1))}$ must be included in $\mathbf{V}_{j}^{(t)}$ and controlled for when estimating unit $j$’s propensity score at time $t$. This adjustment is straightforward when $p$ is known and fixed, and it does not affect our theoretical results. We leave the case where $p$ is unknown or grows with $N$ to future work.
The second relaxation concerns the dependence in $\mathbf{A}_{\mathcal{N}}^{(1:T)}$ across units. Assumption 1 implies that $A_{i}^{(t)}\perp A_{j}^{(t)}\mid\mathbf{V}_{\mathcal{N}}^{(t)}$ for any $i$, $j$, and $t$. However, this may not hold under assignment mechanisms like complete randomization, where treatment allocation is interdependent. In the Supplementary Material, we follow the derivation in Sävje, Aronow and Hudgens (2021) and show that our theoretical results remain valid as long as the dependence is not too strong.
When Assumption 3 fails, variables from future periods, including treatment statuses and time-varying confounders, must also be included in $\mathbf{V}_{j}^{(t)}$ for each unit $j$ in period $t$. As in the case with treatment diffusion, it suffices to control for these variables in the propensity score, and the theoretical results remain unchanged. Finally, we can replace Assumption 5 with bounded moments for $Y_{i}\left(\mathbf{a}^{(1:T)}\right)$. Then, the asymptotic properties of our estimator still follow from the central limit theorem in Kojevnikov, Marmer and Song (2021), although the conditions for Theorem 1 to hold become less intuitive.
A Longer Period
So far, we have focused on settings where $T$ is fixed as $N$ increases. If we also allow $T$ to grow, the number of possible treatment histories expands exponentially, making it infeasible to summarize all AMRs using a saturated MSM. In such cases, researchers must impose structure on the MSM by selecting summary statistics of the treatment history and modeling their influence on the outcome. In the Supplementary Material, we show that if the MSM is saturated with respect to the chosen statistics, the resulting estimates converge to marginalized causal effects generated by those statistics. The same estimator can be used to fit the MSM, and our theoretical results still hold. Importantly, these causal quantities retain substantive meaning even if the selected statistics do not fully capture the effects of the entire treatment history.
Measurement Error in the Proximity Metric
In practice, the proximity metric $d_{ij}$ may be measured with error. For example, interference might depend on commuting time or cultural similarity between two locations, yet researchers only observe straight-line distance. In social network studies, it is common to approximate the full network using self-reported ties, which often fail to capture all relevant interactions in respondents’ lives. However, such measurement error affects the definition of the estimand—that is, the particular form of spillover effect being studied—rather than its identification or estimation. In the extreme case where $d_{ij}$ is completely unavailable, one can still investigate the direct effect of a unit’s treatment history on its own outcome, accounting for interference from others in the sample. Inference becomes more challenging in such settings, as the HAC variance estimator requires researchers to specify a cutoff value $\bar{d}_{N}$. A practical approach is to assess the robustness of variance estimates across a range of $\bar{d}_{N}$ values or to use the conservative variance estimator proposed by Aronow and Samii (2017), which is built upon Young’s inequality.
6 Application
In this section, we illustrate the application of the method using one empirical example in a spatial setting and the other in a social network setting. Simulation results on the method’s performance are reported in the Supplementary Material, which show that the WLS estimator yields consistent estimates and confidence intervals with desirable coverage rates.
6.1 Public Project and Political Support
We first apply our method to the study in Stokes (2016) on the political consequences of building wind turbines in Ontario, Canada, where the Liberal Party government passed laws since 2003 enabling individuals and corporations to build wind turbines, despite objections from local communities. The study collected election data from 6,186 precincts in Ontario over three elections: 2003, 2007, and 2011, and combined the data with the location of each proposed turbine. Precinct $i$ is considered as treated in year $t$ if a turbine project within its boundary has been proposed before the election. The data exhibit the structure of staggered adoption, where once a unit is treated, it remains treated for the remaining periods. The three treatment histories from 2007 to 2011 ($t=1$ and $t=2$) are $(0,0)$, $(0,1)$, $(1,1)$. There were 6,002 never treated precincts, 53 treated since 2007, and 131 treated since 2011.111111The original analysis combines the difference-in-differences estimator with an exposure mapping, hence its validity hinges on the correct model specification. The outcomes of interest are turnout rate and the Liberal Party’s vote share in 2011.
The propensity scores are estimated by a logistic regression model, which controls for the longitude and latitude of each precinct, as well as its turnout rate and vote share for the Liberal Party in the previous election. We construct doughnuts with a radius of 2 km around the geographic center of each treated and untreated precinct, and fit the MSM in the distance range of 0 km to 16 km. The the spatial HAC variance estimator’s cutoff value, $\bar{d}_{N}$, is set at $\max\{20,2*d\}$ km for both outcomes.
The results are presented in the top panel of Figure 3. On the left side are the estimated causal effects for the turnout rate, and on the right side are those for the Liberal Party’s vote share. The black dots in the figures represent estimates for $\mu((1,1);d)-\mu((0,0);d)$, and the gray dots represent estimates for $\mu((0,1);d)-\mu((0,0);d)$. We use superscripts on $\mu(\cdot)$ to indicate the time at which the outcome is measured. The segments denote the $95\%$ confidence intervals of these estimates. The estimates provide evidence for the presence of spillover effects generated by both histories. Being treated in both $2007$ and $2011$ would increase the turnout rate by $7.7\%$ and decrease the Liberal Party’s vote share by $6.4\%$ in a precinct. The effects are also observed among its neighbors that are $2$-$4$ km away, where the turnout rate would increase by $5.0\%$, and the Liberal Party’s vote share would decrease by 6.0%. In precincts treated only in $2011$, treatment assignment history caused a rise of $3.1\%$ in the turnout rate and a decline of $7.0\%$ in the Liberal Party’s vote share. The spillover effects on its neighbors that are $2$-$4$ km away are of similar magnitude. The effects on the outcome variables become close to zero (and statistically insignificant) among precincts that are more than $6$ km away from the treatment, which justifies the choice of the cutoff values.
Figure 3: Results from Applications
(a) Political Consequences of Proposed Wind Turbines
(b) Impacts of Smoking Cessation on Health Outcomes
Notes: Panel (a) presents results replicating Stokes (2016), while Panel (b) displays findings from the FHS. Black and gray dots represent point estimates for different treatment histories and proximity levels. The vertical segments show the corresponding 95% confidence intervals, computed using the spatial/network HAC variance estimator and standard normal critical values.
6.2 Impacts of Smoking Cessation on Health Outcomes
Our second application examines how quitting smoking affects both an individual’s health and that of their social network neighbors. Prior research has established strong links between smoking and adverse health outcomes (Jain et al., 2016). It is natural to conjecture that reducing exposure to second-hand smoke could yield benefits for those nearby. We use data from the Framingham Heart Study (FHS), focusing on the offspring cohort and variables from the first three exams. The average time gap is 8 years between exams 1 and 2, and 5 years between exams 2 and 3. We treat exam 1 as the initial period ($t=0$) and define treatment history using exams 2 and 3 ($t\in\{1,2\}$). $A_{i}^{(t)}=1$ if the individual reports not smoking at time $t$. Outcomes include body weight, blood pressure, cholesterol level, and incidence of cardiovascular disease (CVD). The first three are measured at exam 3 ($t=2$); CVD incidence is defined by the occurrence of symptoms (e.g., stroke) within 10 years after exam 3.121212We examine whether there is any CVD incidence between exam 4 and exam 6. The attrition rate during this period is 16.4%, therefore the influence of righ-censoring is limited. FHS records social ties across multiple dimensions (e.g., spouse, family, friends, neighbors, etc.). We define two individuals as connected if any type of social tie exists between them and include only ties formed prior to exam 2 to avoid endogeneity in network formation.
We restrict our analysis to smokers in the initial period ($A_{i}^{(0)}=1$) and exclude those with missing data or without any social ties, resulting in a final sample of 930 participants. Among them, 347 continued smoking through exams 2 and 3 ($\mathbf{A}_{i}^{(1:2)}=(0,0)$); 413 quit smoking by exam 2 and remained non-smokers in exam 3 ($\mathbf{A}_{i}^{(1:2)}=(1,1)$); 136 quit smoking only by exam 3 ($\mathbf{A}_{i}^{(1:2)}=(0,1)$); and 34 quit in exam 2 but resumed smoking by exam 3 ($\mathbf{A}_{i}^{(1:2)}=(1,0)$). Following the design in Jain et al. (2016), we estimate propensity scores using a logistic regression model that controls for time-invariant covariates (gender, education, height, initial age), as well as treatment status and time-varying covariates (body weight, blood pressure, cholesterol level, and CVD incidence) from the previous period. We set $\mathcal{D}=\{0,1,2,3,4,5\}$ and focus on the effects of treatment histories $(1,1)$ and $(0,1)$ relative to $(0,0)$. These effects capture the impact of quitting smoking on outcomes of interest at exams 2 and 3 , respectively.
The bottom panel of Figure 3 presents our estimates and confidence intervals for both body weight and CVD incidence. The estimand is $\mu\left((0,1);d\right)-\mu\left((0,0);d\right)$ on the left and $\mu\left((1,1);d\right)-\mu\left((0,0);d\right)$ on the right. The results suggest that smoking cessation leads to a significant increase in body weight, with a larger effect for those who quit smoking only by exam 3 (12 pounds) compared to those who quit by exam 2 (5 pounds). These findings are consistent with Jain et al. (2016). Smoking cessation also leads to a significant reduction in CVD risk, with the two treatment histories yielding similar effects (6% vs. 4%). However, we find little evidence of spillover effects on either body weight or CVD incidence, suggesting that reductions in second-hand smoke may have limited impact on others’ health.
Additional results from both applications are provided in the Supplementary Material, including placebo tests supporting the identification assumptions, estimates based on a propensity score model that accounts for treatment diffusion, and estimated effects on blood pressure and cholesterol levels in the FHS.
7 Conclusion
This paper tackles the challenge of causal inference in longitudinal data with an unknown interference structure. We introduce a novel estimand, the AMR, to capture both direct and spillover effects generated by different treatment histories. This estimand is well-defined whenever a proximity metric is available, requiring only minimal knowledge of the underlying interference structure. It can be summarized by an MSM and identified under a slightly generalized form of sequential exchangeability. We show how to estimate any AMR and MSM parameters using a WLS estimator and establish its large-sample properties, including consistency and asymptotic normality, under mild restrictions on the dependence across units. We also develop HAC variance estimators for spillover effect estimates that are ensured to be asymptotically conservative. Our approach bridges methods in longitudinal data analysis and causal inference under interference, providing researchers with a powerful tool to study a broad range of causal effects in spatial and network settings, especially when interactions among units are poorly understood. We illustrate the method’s utility through simulations and two empirical applications, and conclude by discussing possible extensions.
Compared to alternative approaches based on outcome models or exposure mappings, our method avoids risks of model misspecification and the burden of computing exposure probabilities in settings with complex and unknown interference. However, it relies on researchers’ substantive understanding of the treatment assignment mechanism so that the propensity score is either known or can be consistently estimated. Future research should investigate more flexible strategies for propensity score estimation, such as sieve estimators (Hirano, Imbens and Ridder, 2003) or highly adaptive LASSO (Ertefaie, Hejazi and van der Laan, 2020). Another promising direction is to combine our approach with outcome modeling to improve robustness and efficiency.
Supplementary Material
Appendix A Proofs
A.1 Structural Equations under the Identification Assumptions
Under Assumptions 1-3 in the main text, the relationships among variables in our framework can be represented by the following structural equations:
$\mathbf{L}_{i}^{(t)}=f_{\mathbf{L}}\left(\mathbf{A}_{\mathcal{N}_{\mathcal{G}_{N}}(i)}^{1:(t-1)},\mathbf{L}_{\mathcal{N}_{\mathcal{G}_{N}}(i)}^{1:(t-1)},\varepsilon_{\mathbf{L}i}^{(t)}\right),$
$A_{i}^{(t)}=f_{A}\left(\mathbf{A}_{i}^{(1:(t-1))},\mathbf{L}_{i}^{(1:t)},\varepsilon_{{A}i}^{(t)}\right),$
$Y_{i}=f_{Y}\left(\mathbf{A}_{\mathcal{N}_{\mathcal{G}_{N}}(i)}^{(1:T)},\mathbf{L}_{\mathcal{N}_{\mathcal{G}_{N}}(i)}^{(1:T)},\varepsilon_{Yi}\right),$
$\varepsilon_{Ai}^{(t)}\perp\varepsilon_{Aj}^{(t)},\text{ and }\left(\varepsilon_{Yi},\varepsilon_{\mathbf{L}i}^{(t+1)},\mathbf{A}_{i}^{(1:T)},\mathbf{L}_{i}^{(1:T)}\right)\perp\varepsilon_{Aj}^{(t)}\mid\left(\mathbf{A}_{j}^{(1:(t-1))},\mathbf{L}_{j}^{(1:t)}\right)\text{ for any }i,j,t.$
Here, $\varepsilon_{\mathbf{L}i}^{(t)}$, $\varepsilon_{Ai}^{(t)}$, and $\varepsilon_{Yi}$ represent other factors that may influence $\mathbf{L}_{i}^{(t)}$, $A_{i}^{(t)}$, and $Y_{i}$, respectively. They may include a unit’s own attributes, characteristics of its neighbors, and measurement errors. We do not assume that these error terms are identically distributed. $\mathcal{N}_{\mathcal{G}_{N}}(i)$ in the subscripts denotes a unit-specific and unknown neighborhood for unit $i$ in $\mathcal{G}_{N}$. The equations are compatible with the DAG in Figure 1 of the main text and closely resemble those in Ogburn et al. (2020). However, our framework allows for arbitrary heterogeneity in how units respond to treatment histories and does not impose a predefined exposure mapping that summarizes the effects of $\mathbf{A}_{\mathcal{N}}^{(1:T)}$. These equations formally define what unknown interference means and reflect a more realistic setting for spatial or network data, where both $Y_{i}$ and $\mathbf{L}_{i}^{(t)}$ may be shaped by unobserved and idiosyncratic factors, such as unit-specific neighborhood structures. This flexibility introduces the additional challenges for inference discussed below, similar to those encountered in design-based frameworks, although we allow for stochastic variation beyond the treatment assignment mechanism.
A.2 Proof for Proposition 1
For any units $i$ and $j$, we know that
${\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\}Y_{i}}{e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\right)}\right]={\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\}Y_{i}\left(\mathbf{A}_{\mathcal{N}}^{(1:T)}\right)}{e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\right)}\right]$
$=$ ${\rm E}\,\left[{\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\}Y_{i}\left(\mathbf{A}_{\mathcal{N}}^{(1:T)}\right)}{e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\right)}\mid\mathbf{V}_{j}^{(T)}\right]\right]$
$=$ ${\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:(T-1))}=\mathbf{a}^{(1:(T-1))}\}}{e\left(\mathbf{A}_{j}^{(1:(T-1))}=\mathbf{a}^{(1:(T-1))}\right)}{\rm E}\,\left[\frac{\mathbf{1}\{A_{j}^{(T)}=a^{(T)}\}Y_{i}\left(\mathbf{A}_{\mathcal{N}}^{(1:T)}\right)}{P\left(A_{j}^{(T)}=a^{(T)}\mid\mathbf{V}_{j}^{(T)}\right)}\mid\mathbf{V}_{j}^{(T)}\right]\right]$
$=$ ${\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:(T-1))}=\mathbf{a}^{(1:(T-1))}\}}{e\left(\mathbf{A}_{j}^{(1:(T-1))}=\mathbf{a}^{(1:(T-1))}\right)}{\rm E}\,\left[Y_{i}\left(a^{(T)},\mathbf{A}_{j}^{(1:(T-1))},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)\mid\mathbf{V}_{j}^{(T)},A_{j}^{(T)}=a^{(T)}\right]\right]$
$=$ ${\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:(T-1))}=\mathbf{a}^{(1:(T-1))}\}}{e\left(\mathbf{A}_{j}^{(1:(T-1))}=\mathbf{a}^{(1:(T-1))}\right)}{\rm E}\,\left[Y_{i}\left(a^{(T)},\mathbf{A}_{j}^{(1:(T-1))},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)\mid\mathbf{V}_{j}^{(T)}\right]\right]$
$=$ ${\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:(T-1))}=\mathbf{a}^{(1:(T-1))}\}Y_{i}\left(a^{(T)},\mathbf{A}_{j}^{(1:(T-1))},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)}{e\left(\mathbf{A}_{j}^{(1:(T-1))}=\mathbf{a}^{(1:(T-1))}\right)}\right]=\dots$
$=$ ${\rm E}\,\left[Y_{i}\left(\mathbf{a}^{(1:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)\right]=Y_{i;j}\left(\mathbf{a}^{(1:T)}\right).$
The $\dots$ part iterates the same step from period $T-1$ to period $1$. The first equality uses the definition of the potential outcome, while the fourth equality uses the law of total expectation. The fifth equality holds because of sequential exchaneagibility. Therefore, for any $d\in\mathcal{D}$,
$\frac{1}{N}\sum_{j=1}^{N}{\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\}\mu_{j}(d)}{e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\right)}\right]=\frac{1}{N}\sum_{j=1}^{N}{\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\}\mu\left(\big{\{}Y_{i}\big{\}}_{i\in\mathcal{N}};d\right)}{e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\right)}\right]$
$=$ $\frac{1}{N}\sum_{j=1}^{N}\frac{1}{|\Omega_{j}(d)|}\sum_{i=1}^{N}\mathbf{1}\{i\in\Omega_{j}(d)\}{\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\}Y_{i}}{e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\right)}\right]$
$=$ $\frac{1}{N}\sum_{j=1}^{N}\frac{\sum_{i=1}^{N}\mathbf{1}\{i\in\Omega_{j}(d)\}Y_{i;j}\left(\mathbf{a}^{(1:T)}\right)}{|\Omega_{j}(d)|}=\mu\left(\mathbf{a}^{(1:T)};d\right).$
In a saturated MSM, each of the parameters is a linear combination of the AMRs thus can also be identified from data.
A.3 Proof for Theorem 1
We start from the case where $e\left(\mathbf{A}_{i}^{(1:T)}=\mathbf{A}^{(1:T)}\right)$ is known. Consider the first term in the expression of $\hat{\beta}(d)$, namely, $\frac{1}{N}\sum_{i=1}^{N}w_{i}\mathbf{m}\left(\mathbf{A}_{i}^{(1:T)}\right)\mathbf{m}^{\prime}\left(\mathbf{A}_{i}^{(1:T)}\right)$. We can establish the following lemma:
Lemma A.1
Under Assumption 1,
${\rm E}\,\left[\frac{f\left(\mathbf{A}_{i}^{(1:T)}\right)f\left(\mathbf{A}_{j}^{(1:T)}\right)}{e\left(\mathbf{A}_{i}^{(1:T)}=\mathbf{A}^{(1:T)}\right)e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{A}^{(1:T)}\right)}\right]={\rm E}\,\left[\frac{f\left(\mathbf{A}_{i}^{(1:T)}\right)}{e\left(\mathbf{A}_{i}^{(1:T)}=\mathbf{A}^{(1:T)}\right)}\right]{\rm E}\,\left[\frac{f\left(\mathbf{A}_{j}^{(1:T)}\right)}{e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{A}^{(1:T)}\right)}\right]$
for any function $f(\cdot)$ and units $i,j\in\mathcal{N}_{N}$ and $i\neq j$.
Proof: Let $\mathcal{A}^{(1:T)}$ denote all possible values that $\mathbf{A}_{i}^{(1:T)}$ can take, then
${\rm E}\,\left[\frac{f\left(\mathbf{A}_{i}^{(1:T)}\right)}{e\left(\mathbf{A}_{i}^{(1:T)}=\mathbf{A}^{(1:T)}\right)}\right]={\rm E}\,\left[{\rm E}\,\left[\frac{f\left(\mathbf{A}_{i}^{(1:T)}\right)}{e\left(\mathbf{A}_{i}^{(1:T)}=\mathbf{A}^{(1:T)}\right)}\mid\mathbf{V}_{i}^{(T)}\right]\right]$
$=$ ${\rm E}\,\left[\frac{{\rm E}\,\left[\frac{f\left(\mathbf{A}_{i}^{(1:T)}\right)}{P\left(A_{i}^{(T)}=A^{(T)}|\mathbf{V}_{i}^{(T)}\right)}\mid\mathbf{V}_{i}^{(T)},\right]}{\prod_{s=1}^{T-1}P\left(A_{i}^{(s)}=A^{(s)}|\mathbf{V}_{j}^{(s)}\right)}\right]$
$=$ ${\rm E}\,\left[\frac{\sum_{a_{i}^{(T)}=0}^{1}{\rm E}\,\left[\frac{f\left(a_{i}^{(T)},\mathbf{A}_{i}^{(1:(T-1))}\right)}{P\left(A_{i}^{(T)}=a_{i}^{(T)}|\mathbf{V}_{i}^{(T)}\right)}\mid\mathbf{V}_{i}^{(T)},A_{i}^{(T)}=a_{i}^{(T)}\right]P\left(A_{i}^{(T)}=a_{i}^{(T)}\mid\mathbf{V}_{i}^{(T)}\right)}{\prod_{s=1}^{T-1}P\left(A_{i}^{(s)}|\mathbf{V}_{j}^{(s)}\right)}\right]$
$=$ $\sum_{a_{i}^{(T)}=0}^{1}{\rm E}\,\left[\frac{f\left(a_{i}^{(T)},\mathbf{A}_{i}^{1:(T-1)}\right)}{\prod_{s=1}^{T-1}P\left(A_{i}^{(s)}|\mathbf{V}_{j}^{(s)}\right)}\right]=\dots=\sum_{\mathbf{a}_{i}^{(1:T)}\in\mathcal{A}^{(1:T)}}f\left(\mathbf{a}_{i}^{(1:T)}\right).$
Assumption 1 implies that $A_{i}^{(T)}\perp A_{j}^{(T)}\mid\left(\mathbf{V}_{i}^{(T)},\mathbf{V}_{j}^{(T)}\right)$ for any $t$ and $s$ by the rule of weak union. Therefore, for any $a_{i}^{(T)},a_{j}^{(T)}\in\{0,1\}$, $P\left(A_{i}^{(T)}=a_{i}^{(T)},A_{j}^{(T)}=a_{j}^{(T)}\mid\mathbf{V}_{i}^{(T)},\mathbf{V}_{j}^{(T)}\right)=P(A_{i}^{(T)}=a_{i}^{(T)}|\mathbf{V}_{i}^{(T)})P(A_{j}^{(T)}=a_{j}^{(T)}|\mathbf{V}_{j}^{(T)})$, and
${\rm E}\,\left[\frac{f\left(\mathbf{A}_{i}^{(1:T)}\right)f\left(\mathbf{A}_{j}^{(1:T)}\right)}{e\left(\mathbf{A}_{i}^{(1:T)}=\mathbf{A}^{(1:T)}\right)e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{A}^{(1:T)}\right)}\right]$
$=$ ${\rm E}\,\left[{\rm E}\,\left[\frac{f\left(\mathbf{A}_{i}^{(1:T)}\right)f\left(\mathbf{A}_{j}^{(1:T)}\right)}{e\left(\mathbf{A}_{i}^{(1:T)}=\mathbf{A}^{(1:T)}\right)e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{A}^{(1:T)}\right)}\mid\mathbf{V}_{i}^{(T)},\mathbf{V}_{j}^{(T)}\right]\right]$
$=$ ${\rm E}\,\left[\frac{{\rm E}\,\left[\frac{1}{P\left(A_{i}^{(T)}|\mathbf{V}_{i}^{(T)}\right)P\left(A_{j}^{(T)}|\mathbf{V}_{j}^{(T)}\right)}f\left(\mathbf{A}_{i}^{(1:T)}\right)f\left(\mathbf{A}_{j}^{(1:T)}\right)\mid\mathbf{V}_{i}^{(T)},\mathbf{V}_{j}^{(T)}\right]}{\prod_{s=1}^{T-1}P\left(A_{i}^{(s)}|\mathbf{V}_{j}^{(s)}\right)\prod_{s=1}^{T-1}P\left(A_{j}^{(s)}|\mathbf{V}_{j}^{(s)}\right)}\right]$
$=$ ${\rm E}\,\left[\frac{\sum_{a_{i}^{(T)}=0}^{1}\sum_{a_{j}^{(T)}=0}^{1}{\rm E}\,\left[\frac{f\left(a_{i}^{(T)},\mathbf{A}_{i}^{(1:(T-1))}\right)f\left(a_{j}^{(T)},\mathbf{A}_{j}^{(1:(T-1))}\right)}{P\left(A_{i}^{(T)}=a_{i}^{(T)}|\mathbf{V}_{i}^{(T)}\right)P\left(A_{j}^{(T)}=a_{j}^{(T)}|\mathbf{V}_{j}^{(T)}\right)}\mid\mathbf{V}_{i}^{(T)},\mathbf{V}_{j}^{(T)},A_{i}^{(T)}=a_{i}^{(T)},A_{j}^{(T)}=a_{j}^{(T)}\right]}{\frac{1}{P\left(A_{i}^{(T)}=a_{i}^{(T)},A_{j}^{(T)}=a_{j}^{(T)}\mid\mathbf{V}_{i}^{(T)},\mathbf{V}_{j}^{(T)}\right)}\prod_{s=1}^{T-1}P\left(A_{i}^{(s)}|\mathbf{V}_{j}^{(s)}\right)\prod_{s=1}^{T-1}P\left(A_{j}^{(s)}|\mathbf{V}_{j}^{(s)}\right)}\right]$
$=$ $\sum_{a_{i}^{(T)}=0}^{1}\sum_{a_{j}^{(T)}=0}^{1}{\rm E}\,\left[\frac{f\left(a_{i}^{(T)},\mathbf{A}_{i}^{(1:(T-1))}\right)f\left(a_{j}^{(T)},\mathbf{A}_{j}^{(1:(T-1))}\right)}{\prod_{s=1}^{T-1}P\left(A_{i}^{(s)}|\mathbf{V}_{j}^{(s)}\right)\prod_{s=1}^{T-1}P\left(A_{j}^{(s)}|\mathbf{V}_{j}^{(s)}\right)}\right]=\dots$
$=$ $\sum_{\mathbf{a}_{i}^{(1:T)}\in\mathcal{A}^{(1:T)}}\sum_{\mathbf{a}_{j}^{(1:T)}\in\mathcal{A}^{(1:T)}}f\left(\mathbf{a}_{i}^{(1:T)}\right)f\left(\mathbf{a}_{j}^{(1:T)}\right)={\rm E}\,\left[\frac{f\left(\mathbf{A}_{i}^{(1:T)}\right)}{e\left(\mathbf{A}_{i}^{(1:T)}=\mathbf{A}^{(1:T)}\right)}\right]{\rm E}\,\left[\frac{f\left(\mathbf{A}_{j}^{(1:T)}\right)}{e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{A}^{(1:T)}\right)}\right].$
In our framework, $\mathbf{L}_{i}^{(T)}$ and $\mathbf{L}_{j}^{(T)}$ are dependent because both are functions of $\mathbf{A}_{\mathcal{N}}^{(1:T)}$. As a result, for any function $f(\cdot)$, $f\left(\mathbf{A}_{i}^{(1:T)}\right)$ and $f\left(\mathbf{A}_{j}^{(1:T)}\right)$ are generally dependent as well. Nevertheless, Lemma A.1 suggests that once normalized by the weights, $\frac{1}{e\left(\mathbf{A}_{i}^{(1:T)}=\mathbf{A}^{(1:T)}\right)}$ and $\frac{1}{e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{A}^{(1:T)}\right)}$, they become uncorrelated. Due to positivity and the binary nature of $A_{i}^{(T)}$, $w_{i}\mathbf{m}\left(\mathbf{A}_{i}^{(1:T)}\right)\mathbf{m}^{\prime}\left(\mathbf{A}_{i}^{(1:T)}\right)$ is uniformly bounded. Then, using Lemma A.1, we can see that
${\rm Var}\,\left[\frac{1}{N}\sum_{i=1}^{N}w_{i}\mathbf{m}\left(\mathbf{A}_{i}^{(1:T)}\right)\mathbf{m}^{\prime}\left(\mathbf{A}_{i}^{(1:T)}\right)\right]=\frac{1}{N^{2}}\sum_{i=1}^{N}{\rm Var}\,\left[w_{i}\mathbf{m}\left(\mathbf{A}_{i}^{(1:T)}\right)\mathbf{m}^{\prime}\left(\mathbf{A}_{i}^{(1:T)}\right)\right]\to 0,$
as $N\to\infty$. Therefore, the term $\frac{1}{N}\sum_{i=1}^{N}w_{i}\mathbf{m}\left(\mathbf{A}_{i}^{(1:T)}\right)\mathbf{m}^{\prime}\left(\mathbf{A}_{i}^{(1:T)}\right)$ is consistent for its expectation, $\frac{1}{N}\sum_{i=1}^{N}{\rm E}\,\left[w_{i}\mathbf{m}\left(\mathbf{A}_{i}^{(1:T)}\right)\mathbf{m}^{\prime}\left(\mathbf{A}_{i}^{(1:T)}\right)\right]$. We use $\Sigma_{A}$ to represent the normalized variance $\frac{1}{N}\sum_{i=1}^{N}{\rm Var}\,\left[w_{i}\mathbf{m}\left(\mathbf{A}_{i}^{(1:T)}\right)\mathbf{m}^{\prime}\left(\mathbf{A}_{i}^{(1:T)}\right)\right]$.
Next, we turn to the second term, $\frac{1}{N}\sum_{i=1}^{N}\mathbf{Z}_{i}(d)Y_{i}$. Let’s define $U_{i}(d)=\kappa^{\prime}\mathbf{Z}_{i}(d)Y_{i}$, where $\kappa$ is a vector of weights that sum up to $1$. We first show that the sequence $\{U_{i}(d)\}_{i=1}^{N}$ satisfies a property termed as $\psi$-dependence by Kojevnikov, Marmer and Song (2021). For any subset $\mathcal{H}\subset\mathcal{N}_{N}$ with $|\mathcal{H}|=h$, we denote the collection of bounded Lipchitz functions on it with the associated Lipchitz constant $L_{h}$ as $\mathcal{L}_{h}$. Consider an independent copy of the entire treatment history, $\left(\mathbf{A}^{(1:T)}_{\mathcal{N}_{N}}\right)^{\dagger}$. Since both $\kappa^{\prime}\mathbf{Z}_{i}(d)$ and $Y_{i}$ are functions of the entire treatment history, we can define $U_{i}^{(p,\dagger)}(d)$ as the value of $U_{i}(d)$ under $\left(\mathbf{A}^{(1:T)}_{\mathcal{N}_{\mathcal{G}_{N}}(i,p)},\left(\mathbf{A}^{(1:T)}_{\mathcal{N}_{N}\setminus\mathcal{N}_{\mathcal{G}_{N}}(i,p)}\right)^{\dagger}\right)$. $U_{i}^{(p,\ddagger)}(d)$ is similarly defined under a different copy, $\left(\mathbf{A}^{(1:T)}_{\mathcal{N}_{N}}\right)^{\ddagger}$.
We further define $\mathbf{U}_{\mathcal{H}}(d)=(\dots,U_{i}(d),\dots)^{\prime}_{i\in\mathcal{H}}$ and $\mathbf{U}_{\mathcal{H}}^{(p,\dagger)}(d)$ similarly. For any subset ~$\tilde{\mathcal{H}}\subset\mathcal{N}_{N}$ such that $\min_{i\in\mathcal{H},j\in\tilde{\mathcal{H}}}d_{ij}\geq p$ and $\mathbf{U}_{\tilde{\mathcal{H}}}^{(p,\ddagger)}(d)$ defined on it, it is straightforward to see that ${\rm Cov}\,\left[f_{h}\left(\mathbf{U}_{\mathcal{H}}^{(\lfloor\frac{p}{2}\rfloor,\dagger)}(d)\right),f_{\tilde{h}}\left(\mathbf{U}_{\tilde{\mathcal{H}}}^{(\lfloor\frac{p}{2}\rfloor,\ddagger)}(d)\right)\right]=0$ if $p>2*d$. Then, for any $f_{h}\in\mathcal{L}_{h}$ and $f_{\tilde{h}}\in\mathcal{L}_{\tilde{h}}$, we know that $\Bigg{|}{\rm Cov}\,\left[f_{h}\left(\mathbf{U}_{\mathcal{H}}(d)\right),f_{\tilde{h}}\left(\mathbf{U}_{\tilde{\mathcal{H}}}(d)\right)\right]\Bigg{|}\leq 2||f_{h}||_{\infty}||f_{\tilde{h}}||_{\infty}$ for $p\leq 2*d$, and for $p>2*d$,
$\Bigg{|}{\rm Cov}\,\left[f_{h}\left(\mathbf{U}_{\mathcal{H}}(d)\right),f_{\tilde{h}}\left(\mathbf{U}_{\tilde{\mathcal{H}}}(d)\right)\right]\Bigg{|}$
$=$ $\Bigg{|}{\rm Cov}\,\left[f_{h}\left(\mathbf{U}_{\mathcal{H}}(d)\right),f_{\tilde{h}}\left(\mathbf{U}_{\tilde{\mathcal{H}}}(d)\right)\right]-{\rm Cov}\,\left[f_{h}\left(\mathbf{U}_{\mathcal{H}}^{(\lfloor\frac{p}{2}\rfloor,\dagger)}(d)\right),f_{\tilde{h}}\left(\mathbf{U}_{\tilde{\mathcal{H}}}^{(\lfloor\frac{p}{2}\rfloor,\ddagger)}(d)\right)\right]\Bigg{|}$
$=$ $\Bigg{|}{\rm Cov}\,\left[f_{h}\left(\mathbf{U}_{\mathcal{H}}(d)\right),f_{\tilde{h}}\left(\mathbf{U}_{\tilde{\mathcal{H}}}(d)\right)\right]-{\rm Cov}\,\left[f_{h}\left(\mathbf{U}_{\mathcal{H}}^{(\lfloor\frac{p}{2}\rfloor,\dagger)}(d)\right),f_{\tilde{h}}\left(\mathbf{U}_{\tilde{\mathcal{H}}}(d)\right)\right]$
$+{\rm Cov}\,\left[f_{h}\left(\mathbf{U}_{\mathcal{H}}^{(\lfloor\frac{p}{2}\rfloor,\dagger)}(d)\right),f_{\tilde{h}}\left(\mathbf{U}_{\tilde{\mathcal{H}}}(d)\right)\right]-{\rm Cov}\,\left[f_{h}\left(\mathbf{U}_{\mathcal{H}}^{(\lfloor\frac{p}{2}\rfloor,\dagger)}(d)\right),f_{\tilde{h}}\left(\mathbf{U}_{\tilde{\mathcal{H}}}^{(\lfloor\frac{p}{2}\rfloor,\ddagger)}\right)(d)\right]\Bigg{|}$
$\leq$ $\Bigg{|}{\rm Cov}\,\left[f_{h}\left(\mathbf{U}_{\mathcal{H}}(d)\right)-f_{h}\left(\mathbf{U}_{\mathcal{H}}^{(\lfloor\frac{p}{2}\rfloor,\dagger)}(d)\right),f_{\tilde{h}}\left(\mathbf{U}_{\tilde{\mathcal{H}}}(d)\right)\right]\Bigg{|}$
$+\Bigg{|}{\rm Cov}\,\left[f_{h}\left(\mathbf{U}_{\mathcal{H}}^{(\lfloor\frac{p}{2}\rfloor,\dagger)}(d)\right),f_{\tilde{h}}\left(\mathbf{U}_{\tilde{\mathcal{H}}}(d)\right)-f_{\tilde{h}}\left(\mathbf{U}_{\tilde{\mathcal{H}}}^{(\lfloor\frac{p}{2}\rfloor,\ddagger)}(d)\right)\right]\Bigg{|}$
$\leq$ $2\Big{|}||f_{\tilde{h}}||_{\infty}{\rm E}\,\left[f_{h}\left(\mathbf{U}_{\mathcal{H}}(d)\right)-f_{h}\left(\mathbf{U}_{\mathcal{H}}^{(\lfloor\frac{p}{2}\rfloor,\dagger)}(d)\right)\right]\Big{|}+2\Big{|}||f_{h}||_{\infty}{\rm E}\,\left[f_{\tilde{h}}\left(\mathbf{U}_{\tilde{\mathcal{H}}}(d)\right)-f_{\tilde{h}}\left(\mathbf{U}_{\tilde{\mathcal{H}}}^{(\lfloor\frac{p}{2}\rfloor,\ddagger)}(d)\right)\right]\Big{|}$
$\leq$ $2||f_{\tilde{h}}||_{\infty}hL_{h}\theta_{N,\lfloor\frac{p}{2}\rfloor}(d)+2||f_{h}||_{\infty}\tilde{h}L_{\tilde{h}}\theta_{N,\lfloor\frac{p}{2}\rfloor}(d).$
As $Y_{i}$ is uniformly bounded, so is $U_{i}$. Then, Assumption 3.3 in Kojevnikov, Marmer and Song (2021) is satisfied. It is straightforward to verify that conditions in Theorem 1 imply Assumption 3.4 in Kojevnikov, Marmer and Song (2021). The asymptotic normality of $\frac{1}{N}\sum_{i=1}^{N}U_{i}$ follows from Theorem 3.2 in Kojevnikov, Marmer and Song (2021). Using the Cramer-Wold device, we know that $\frac{1}{N}\sum_{i=1}^{N}\mathbf{Z}_{i}(d)Y_{i}$ is also asymptotically normal:
$\frac{\left(\Sigma_{Y}(d)\right)^{-\frac{1}{2}}}{\sqrt{N}}\sum_{i=1}^{N}\left[\mathbf{Z}_{i}(d)Y_{i}-{\rm E}\,\left[\mathbf{Z}_{i}(d)Y_{i}\right]\right]\to\mathcal{N}(0,\mathbf{I}),$
Theorem 1 can then be derived from Slutsky’s Theorem, with $\Sigma_{N}(d)=\left(\Sigma_{A}\right)^{-1}\Sigma_{Y}(d)\left(\Sigma_{A}\right)^{-1}$.
Consider our simulated example in the main text with $t\in\{0,1,2\}$ and $\mathbf{m}\left(\mathbf{A}^{(1:T)}\right)=\left(1,A_{1},A_{2},A_{1}A_{2}\right)^{\prime}$. Then,
$\left(\frac{1}{N}\sum_{i=1}^{N}w_{i}\mathbf{m}\left(\mathbf{A}_{i}^{(1:T)}\right)\mathbf{m}^{\prime}\left(\mathbf{A}_{i}^{(1:T)}\right)\right)^{-1}\to\begin{pmatrix}1&-1&-1&1\\ -1&2&1&-2\\ -1&1&2&-2\\ 1&-2&-2&4\end{pmatrix},$
while
$\frac{1}{N}\sum_{i=1}^{N}\mathbf{Z}_{i}(d)Y_{i}\to\begin{pmatrix}\mu\left((0,0);d\right)+\mu\left((0,1);d\right)+\mu\left((1,0);d\right)+\mu\left((1,1);d\right)\\ \mu\left((1,0);d\right)+\mu\left((1,1);d\right)\\ \mu\left((0,1);d\right)+\mu\left((1,1);d\right)\\ \mu\left((1,1);d\right)\end{pmatrix}.$
Therefore, $\hat{\beta}(d)\to\beta(d)=\left(\frac{1}{N}\sum_{i=1}^{N}{\rm E}\,\left[w_{i}\mathbf{m}\left(\mathbf{A}_{i}^{(1:T)}\right)\mathbf{m}^{\prime}\left(\mathbf{A}_{i}^{(1:T)}\right)\right]\right)^{-1}\frac{1}{N}\sum_{i=1}^{N}{\rm E}\,\left[\mathbf{Z}_{i}(d)Y_{i}\right]$, where
$\beta(d)=\begin{pmatrix}\mu\left((0,0);d\right)\\ \mu\left((0,1);d\right)-\mu\left((0,0);d\right)\\ \mu\left((1,0);d\right)-\mu\left((0,0);d\right)\\ \mu\left((1,1);d\right)-\mu\left((0,1);d\right)-\mu\left((1,0);d\right)+\mu\left((0,0);d\right)\end{pmatrix}.$
When the propensity score needs to be estimated from data, there is an additional part in the variance, driven by the uncertainty from estimating $e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{A}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)$. Its expression can be derived using the standard M-estimation theory, as in Lunceford and Davidian (2004). We present the expression of the additional part when the propensity score is estimated via a logistic regression:
$P\left(A_{j}^{(t)}=1\mid\mathbf{V}_{j}^{(t)}\right)=\frac{\exp\left(\mathbf{V}_{j}^{(t)^{\prime}}\gamma_{t}\right)}{1+\exp\left(\mathbf{V}_{j}^{(t)^{\prime}}\gamma_{t}\right)}=g\left(\mathbf{V}_{j}^{(t)};\gamma_{t}\right).$
We know that $\frac{}{}$ $\Sigma_{\gamma_{t}}={\rm Var}\,\left[\hat{\gamma}_{t}\right]=\sum_{j=1}^{N}\mathbf{V}_{j}^{(t)^{\prime}}g\left(\mathbf{V}_{j}^{(t)};\gamma_{t}\right)\left(1-g\left(\mathbf{V}_{j}^{(t)};\gamma_{t}\right)\right)\mathbf{V}_{j}^{(t)}$. Remember that the regression residual for unit $j$ from the WLS is denoted as $\hat{\varepsilon}_{j}$, and define $w_{j}^{SW}=\prod_{s=1}^{T}P\left(A_{j}^{(s)}=A^{(s)}\mid\mathbf{A}_{j}^{(1:(s-1))}=\mathbf{A}^{(1:(s-1))}\right)$. Further define $\mathbf{H}_{t}(d)$ as
$\sum_{j=1}^{N}\frac{w_{j}^{SW}}{\prod_{s\neq t}g\left(\mathbf{V}_{j}^{(s)};\gamma_{s}\right)}\left(\frac{\left(1-g\left(\mathbf{V}_{j}^{(t)};\gamma_{t}\right)\right)A_{j}^{(t)}}{g\left(\mathbf{V}_{j}^{(t)};\gamma_{t}\right)}+\frac{g\left(\mathbf{V}_{j}^{(t)};\gamma_{t}\right)\left(1-A_{j}^{(t)}\right)}{1-g\left(\mathbf{V}_{j}^{(t)};\gamma_{t}\right)}\right)\hat{\varepsilon}_{j}\mathbf{V}_{j}^{(t)^{\prime}}\mathbf{m}\left(\mathbf{A}_{j}^{(1:T)}\right).$
Then, the additional part in the variance equals
$\Sigma_{adj}(d)=-\sum_{t=1}^{T}\mathbf{H}_{t}^{\prime}(d)\Sigma_{\gamma_{t}}\mathbf{H}_{t}(d).$
The final variance equals $\Sigma_{N}(d)+\Sigma_{adj}(d)$.
A.4 Proof for Theorem 2
First, note that $\frac{\mathbf{M}^{\prime}\mathbf{W}\mathbf{M}}{N}\to\Sigma_{A}$ as $N\to\infty$ by Lemma A.1 and standard theory. Define $\mathbf{Q}^{\prime}_{i}(d)=\sum_{k=1}^{N}\frac{\mathbf{1}\{i\in\Omega_{k}(d)\}w_{k}\mathbf{m}\left(\mathbf{A}_{k}^{(1:T)}\right)\mathbf{m}^{\prime}\left(\mathbf{A}_{k}^{(1:T)}\right)}{|\Omega_{k}(d)|}$. Then, $\hat{\boldsymbol{\varepsilon}}_{Yi}(d)=\mathbf{Z}_{i}(d)Y_{i}-\mathbf{Q}^{\prime}_{i}(d)\hat{\boldsymbol{\beta}}(d)$. It is straightforward to verify that $\frac{1}{N}\sum_{i=1}^{N}\mathbf{Q}^{\prime}_{i}(d)=\frac{1}{N}\sum_{i=1}^{N}w_{i}\mathbf{m}\left(\mathbf{A}_{i}^{(1:T)}\right)\mathbf{m}^{\prime}\left(\mathbf{A}_{i}^{(1:T)}\right)$, which implies that $\boldsymbol{\beta}(d)=\left(\frac{1}{N}\sum_{i=1}^{N}{\rm E}\,\left[\mathbf{Q}^{\prime}_{i}(d)\right]\right)^{-1}\left(\frac{1}{N}\sum_{i=1}^{N}{\rm E}\,\left[\mathbf{Z}_{i}(d)Y_{i}\right]\right)$ and $\hat{\boldsymbol{\beta}}(d)=\left(\frac{1}{N}\sum_{i=1}^{N}\mathbf{Q}^{\prime}_{i}(d)\right)^{-1}\left(\frac{1}{N}\sum_{i=1}^{N}\mathbf{Z}_{i}(d)Y_{i}\right)$. Next, we prove the following lemma concerning the terms $\{\mathbf{Z}_{i}(d)Y_{i}\}_{i\in\mathcal{N}_{N}}$ and $\{\mathbf{Q}_{i}(d)\}_{i\in\mathcal{N}_{N}}$:
Lemma A.2
Under Assumptions 1-6,
$\frac{1}{N}\sum_{i=1}^{N}\sum_{j:\{d_{ij}\leq\bar{d}_{N}\}}\kappa_{i}\left[\mathbf{Q}_{i}(d)-{\rm E}\,\left[\mathbf{Q}_{i}(d)\right]\right]\to 0,\frac{1}{N}\sum_{i=1}^{N}\sum_{j:\{d_{ij}\leq\bar{d}_{N}\}}\kappa_{i}\left[\mathbf{Z}_{i}(d)Y_{i}-{\rm E}\,\left[\mathbf{Z}_{i}(d)Y_{i}\right]\right]\to 0,$
for any uniformly bounded weights $\{\kappa_{i}\}_{i\in\mathcal{N}_{N}}$ and units $i,j\in\mathcal{N}_{N}$.
Proof: See the proof of Theorem 4 in Leung (2022). Note that ${\rm Cov}\,\left[\mathbf{Q}_{i}(d),\mathbf{Q}^{\prime}_{j}(d)\right]=0$ if $d_{ij}\geq 2*d$.
Using the same derivation as in Section 4 and standard theory on estimating equations, we have
$\Sigma_{Y}(d)=$ $\frac{1}{N}\sum_{i=1}^{N}\sum_{j:\{d_{ij}\leq\bar{d}_{N}\}}{\rm E}\,\left[\left[\boldsymbol{\varepsilon}_{Yi}(d)-{\rm E}\,\left[\boldsymbol{\varepsilon}_{Yi}(d)\right]\right]\left[\boldsymbol{\varepsilon}_{Yj}(d)-{\rm E}\,\left[\boldsymbol{\varepsilon}_{Yj}(d)\right]\right]\right]^{\prime},$
where $\boldsymbol{\varepsilon}_{Yi}(d)=\mathbf{Z}_{i}(d)Y_{i}-\mathbf{Q}^{\prime}_{i}(d)\boldsymbol{\beta}(d)$. Note that ${\rm E}\,\left[\boldsymbol{\varepsilon}_{Yi}(d)\right]\neq 0$ as the outcomes are not identically distributed. Let’s define
$\widehat{\Sigma}_{Y}(d)=\frac{1}{N}\hat{\boldsymbol{\varepsilon}}^{\prime}_{Y}(d)\mathbf{K}_{\bar{d}_{N}}\hat{\boldsymbol{\varepsilon}}_{Y}(d)=\frac{1}{N}\sum_{i=1}^{N}\sum_{j:\{d_{ij}\leq\bar{d}_{N}\}}\hat{\boldsymbol{\varepsilon}}_{Yi}(d)\hat{\boldsymbol{\varepsilon}}^{\prime}_{Yj}(d),$
$\tilde{\Sigma}_{Y}(d)=\frac{1}{N}\sum_{i=1}^{N}\sum_{j:\{d_{ij}\leq\bar{d}_{N}\}}\left[\boldsymbol{\varepsilon}_{Yi}(d)-{\rm E}\,\left[\boldsymbol{\varepsilon}_{Yi}(d)\right]\right]\left[\boldsymbol{\varepsilon}_{Yj}(d)-{\rm E}\,\left[\boldsymbol{\varepsilon}_{Yj}(d)\right]\right]^{\prime},\text{ and }$
$\check{\Sigma}_{Y}(d)=\frac{1}{N}\sum_{i=1}^{N}\sum_{j:\{d_{ij}\leq\bar{d}_{N}\}}\boldsymbol{\varepsilon}_{Yi}(d)\boldsymbol{\varepsilon}^{\prime}_{Yj}(d).$
Kojevnikov, Marmer and Song (2021) show that $\tilde{\Sigma}_{Y}(d)\to\Sigma_{Y}(d)$ as $N\to\infty$. Let $\mathbf{R}(d)=\frac{1}{N}\sum_{i=1}^{N}\sum_{j:\{d_{ij}\leq\bar{d}_{N}\}}{\rm E}\,\left[\boldsymbol{\varepsilon}_{Yi}(d)\right]{\rm E}\,\left[\boldsymbol{\varepsilon}_{Yj}^{\prime}(d)\right]$. Then, by Lemma A.2,
$\check{\Sigma}_{Y}(d)-\tilde{\Sigma}_{Y}(d)$
$=$ $\frac{1}{N}\sum_{i=1}^{N}\sum_{j:\{d_{ij}\leq\bar{d}_{N}\}}\left(\boldsymbol{\varepsilon}_{Yi}(d){\rm E}\,\left[\boldsymbol{\varepsilon}_{Yj}^{\prime}(d)\right]+{\rm E}\,\left[\boldsymbol{\varepsilon}_{Yi}(d)\right]\boldsymbol{\varepsilon}^{\prime}_{Yj}(d)-{\rm E}\,\left[\boldsymbol{\varepsilon}_{Yi}(d)\right]{\rm E}\,\left[\boldsymbol{\varepsilon}_{Yj}^{\prime}(d)\right]\right)$
$=$ $\frac{1}{N}\sum_{i=1}^{N}\sum_{j:\{d_{ij}\leq\bar{d}_{N}\}}\left(\boldsymbol{\varepsilon}_{Yi}(d)-{\rm E}\,\left[\boldsymbol{\varepsilon}_{Yi}(d)\right]\right){\rm E}\,\left[\boldsymbol{\varepsilon}_{Yj}^{\prime}(d)\right]$
$+\frac{1}{N}\sum_{i=1}^{N}\sum_{j:\{d_{ij}\leq\bar{d}_{N}\}}{\rm E}\,\left[\boldsymbol{\varepsilon}_{Yi}(d)\right]\left(\boldsymbol{\varepsilon}_{Yj}(d)-{\rm E}\,\left[\boldsymbol{\varepsilon}_{Yj}(d)\right]\right)^{\prime}+\frac{1}{N}\sum_{i=1}^{N}\sum_{j:\{d_{ij}\leq\bar{d}_{N}\}}{\rm E}\,\left[\boldsymbol{\varepsilon}_{Yi}(d)\right]{\rm E}\,\left[\boldsymbol{\varepsilon}_{Yj}^{\prime}(d)\right]$
$=$ $\frac{1}{N}\sum_{i=1}^{N}\left[\mathbf{Z}_{i}(d)Y_{i}-{\rm E}\,\left[\mathbf{Z}_{i}(d)Y_{i}\right]\right]\left(\sum_{j=1}^{N}\mathbf{1}\{d_{ij}\leq\bar{d}_{N}\}\right){\rm E}\,\left[\boldsymbol{\varepsilon}_{Yj}^{\prime}(d)\right]$
$+\frac{1}{N}\sum_{i=1}^{N}\left(\sum_{j=1}^{N}\mathbf{1}\{d_{ij}\leq\bar{d}_{N}\}\right){\rm E}\,\left[\boldsymbol{\varepsilon}_{Yi}(d)\right]\left[\mathbf{Z}_{j}(d)Y_{j}-{\rm E}\,\left[\mathbf{Z}_{j}(d)Y_{j}\right]\right]^{\prime}$
$+\frac{1}{N}\sum_{i=1}^{N}\left[\mathbf{Q}_{i}(d)-{\rm E}\,\left[\mathbf{Q}_{i}(d)\right]\right]\left(\sum_{j=1}^{N}\mathbf{1}\{d_{ij}\leq\bar{d}_{N}\}\right){\rm E}\,\left[\boldsymbol{\varepsilon}_{Yj}^{\prime}(d)\right]$
$+\frac{1}{N}\sum_{i=1}^{N}\left(\sum_{j=1}^{N}\mathbf{1}\{d_{ij}\leq\bar{d}_{N}\}\right){\rm E}\,\left[\boldsymbol{\varepsilon}_{Yi}(d)\right]\left[\mathbf{Q}_{j}(d)-{\rm E}\,\left[\mathbf{Q}_{j}(d)\right]\right]^{\prime}+\mathbf{R}(d)\to\mathbf{R}(d)$
Moreover,
$\mid\widehat{\Sigma}_{Y}(d)-\check{\Sigma}_{Y}(d)\mid$
$=$ $\Bigg{|}\frac{1}{N}\sum_{i=1}^{N}\sum_{j:\{d_{ij}\leq\bar{d}_{N}\}}\left(\mathbf{Q}^{\prime}_{i}(d)\left(\boldsymbol{\beta}(d)+\hat{\boldsymbol{\beta}}(d)\right)/2-\mathbf{Z}_{i}(d)Y_{i}\right)\left(\hat{\boldsymbol{\beta}}(d)-\boldsymbol{\beta}(d)\right)^{\prime}\mathbf{Q}_{j}(d)$
$+\frac{1}{N}\sum_{i=1}^{N}\sum_{j:\{d_{ij}\leq\bar{d}_{N}\}}\mathbf{Q}^{\prime}_{i}(d)\left(\hat{\boldsymbol{\beta}}(d)-\boldsymbol{\beta}(d)\right)\left(\left(\boldsymbol{\beta}(d)+\hat{\boldsymbol{\beta}}(d)\right)^{\prime}\mathbf{Q}_{j}(d)/2-\mathbf{Z}_{i}(d)Y_{i}\right)\Bigg{|}$
$\leq$ $\frac{1}{N}\sum_{i=1}^{N}\sum_{j:\{d_{ij}\leq\bar{d}_{N}\}}\frac{C}{\sqrt{N}}=\frac{CM_{N}\left(\bar{d}_{N},1\right)}{\sqrt{N}}\to 0.$
Combining these results, we conclude that $\widehat{\Sigma}_{Y}(d)\to\Sigma_{Y}(d)+\mathbf{R}(d)$.
Finally, note that $\widehat{\Sigma}^{+}_{Y}(d)=\widehat{\Sigma}_{Y}(d)+\widehat{\Sigma}^{-}_{Y}(d)$. We can similarly show that
$\widehat{\Sigma}^{-}_{Y}(d)\to\Sigma^{-}_{Y}(d)+\mathbf{R}^{-}(d).$
Therefore,
$\widehat{\Sigma}^{+}_{Y}(d)\to$ $\Sigma_{Y}(d)+\mathbf{R}(d)+\Sigma^{-}_{Y}(d)+\mathbf{R}^{-}(d)$
$=$ $\Sigma_{Y}(d)+\mathbf{R}^{+}(d)+\Sigma^{-}_{Y}(d)\succeq\Sigma_{Y}(d).$
A.5 Technical Details of the Extensions
Placebo Tests
We demonstrate the logic of conducting placebo tests using the DAG in Figure A.1. For simplicity, we only keep paths that are relevant for causal identification. In this toy example, two units, $\{i,j\}$, are observed over three periods, $\{0,1,2\}$. We assume that sequential exchangeability holds conditional on each unit’s treatment status in the previous period and the current value of time-varying confounders. Under this assumption, there is no direct path from $\mathbf{L}_{i}^{(1)}$ to $A_{i}^{(2)}$.
Now consider a time-varying unobserved variable, $U_{i}^{(t)}$, that influences both $A_{i}^{(t)}$ and $\mathbf{L}_{i}^{(t+1)}$, as shown by dashed arrows in the DAG. In this case, sequential exchangeability is violated. A backdoor path arises between $A_{i}^{(2)}$ and $\mathbf{L}_{i}^{(1)}$, even when conditioning on $A_{i}^{(1)}$ and $\mathbf{L}_{i}^{(2)}$. As a result, we would expect to find a spurious effect generated by $A_{i}^{(2)}$ on $\mathbf{L}_{i}^{(1)}$ using the WLS estimator, which can serve as evidence against the identification assumptions.
Figure A.1: A DAG illustration for placebo tests
[…]
Notes: The DAG depicts two units, $\{i,j\}$, observed over three periods, $\{0,1,2\}$. Variables are represented by circles and causal paths by arrows. White circles denote treatment, gray circles represent time-varying confounders, and black circles correspond to outcomes. Black arrows indicate relationships influencing treatment assignment, gray arrows represent other potential dependencies between variables, and dashed arrows show potential influences from an unobservable variable.
Dependence in Treatment Assignment
To characterize the degree of dependence in treatment status between two units in period $t$, we follow Sävje, Aronow and Hudgens (2021) and define the conditional alpha-mixing coefficient of two random variables, $X_{1}$ and $X_{2}$, given covariates $\mathbf{V}$, as
$\alpha(X_{1},X_{2};\mathbf{V})=\sup_{\begin{subarray}{c}x_{1}\in\sigma(X_{1}|\mathbf{V}),\\ x_{2}\in\sigma(X_{2}|\mathbf{V})\end{subarray}}|P(x_{1}\cap x_{2}|\mathbf{V})-P(x_{1}|\mathbf{V})P(x_{2}|\mathbf{V})|.$
Let $\check{\mathbf{A}}_{\mathcal{N}\setminus\{j\}}^{(t)}(d)$ the subset of treatment assignments in period $t$, excluding $A_{j}^{(t)}$, that affect $\mu_{j}(d)$. We use an indicator $k_{ij}$ to represent whether $\mu_{i}(d)$ and $\mu_{j}(d)$ are dependent. The conditional internal average mixing coefficient in period $t$ is then defined as $\alpha_{INT,t}=\sum_{j=1}^{N}\alpha\left(A_{j}^{(t)},\check{\mathbf{A}}_{\mathcal{N}\setminus\{j\}}^{(t)}(d)\mid\mathbf{V}_{\mathcal{N}}^{(t)}\right)$, and the conditional external average mixing coefficient in period $t$ is defined as $\alpha_{EXT,t}=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{N}(1-g_{ij})\alpha\left(\check{\mathbf{A}}_{\mathcal{N}\setminus\{i\}}^{(t)}(d),\check{\mathbf{A}}_{\mathcal{N}\setminus\{j\}}^{(t)}(d)\mid\mathbf{V}_{\mathcal{N}}^{(t)}\right)$. Theorem 1 holds holds under the following assumption:
Assumption A.1 (Design mixing and design separation)
$\alpha_{INT,t}=o(N),\alpha_{EXT,t}=o(N),$
for any period $t$.
This assumption is satisfied by many common treatment assignment mechanisms, such as complete randomization. Formal proofs can be found in Sävje, Aronow and Hudgens (2021).
Marginal Structural Models with a Large $T$
When $T$ is large, we can consider summarizing the effects of any treatment history using a statistic $\mathcal{T}\left(\mathbf{a}^{(1:T)}\right)$. A common choice is the number of periods under treatment: $\mathcal{T}\left(\mathbf{a}^{(1:T)}\right)=\sum_{t=1}^{T}a^{(t)}$. We then define AMRs for each $d$ and each possible value $\mathcal{T}$ of $\mathcal{T}\left(\mathbf{a}^{(1:T)}\right)$:
$Y_{i;j}\left(\mathcal{T}\right)\coloneqq{\rm E}\,\left[Y_{i}\left(\mathcal{T}\left(\mathbf{A}_{j}^{(1:T)}\right)=\mathcal{T};\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)\right],$
$\mu\left(\big{\{}Y_{i;j}\left(\mathcal{T}\right)\big{\}}_{i\in\mathcal{N}};\Omega_{j}(d)\right)\coloneqq\frac{\sum_{i=1}^{N}\mathbf{1}\{i\in\Omega_{j}(d)\}Y_{i;j}\left(\mathcal{T}\right)}{|\Omega_{j}(d)|},$
$\mu\left(\mathcal{T};d\right)\coloneqq\frac{1}{N}\sum_{j=1}^{N}\mu_{j}\left(\big{\{}Y_{i;j}\left(\mathcal{T}\right)\big{\}}_{i\in\mathcal{N}};d\right).$
As $\mu\left(\mathcal{T};d\right)$ only depends on $\mathcal{T}$, we can summarize all its possible values using a saturated MSM:
$\mu\left(\mathcal{T};d\right)=\mathbf{m}\left(\mathcal{T}\right)^{\prime}\beta(d).$
By the same logic, each $\mu\left(\mathcal{T};d\right)$ can be identified under Assumption 1, and we can estimate $\beta(d)$ using the WLS estimator.
Appendix B Extra Evidence from Simulation and Applications
B.1 Examples of $d_{ij}$ and $\Omega_{j}(d)$
Table B.1: Examples of $d_{ij}$ and $\Omega_{j}(d)$
$d_{ij}$ $\Omega_{j}(d)$
Euclidean or geodesic distance Donut: $\{i\in\mathcal{N}:d-\kappa<d_{ij}\leq d\}$
[…]
Travel time or accessibility Disk: $\{i\in\mathcal{N}:d_{ij}\leq d\}$
Shortest path length in a graph $d$th-degree neighbors: $\{i\in\mathcal{N}:d_{ij}=d\}$
Cultural similarity or social distance Conditional set: $\{i\in\mathcal{N}:d_{ij}=d,\mathbf{X}_{i}=\mathbf{x}\}$
Notes: This table lists common choices for the proximity metric $d_{ij}$ (left column) and the set $\Omega_{j}(d)$ (right column). The two columns are not aligned row-by-row: each entry illustrates a standalone option and may be combined with others in practice. $\mathbf{X}_{i}$ in the final row denotes covariates of unit $i$, which allows researchers to define neighborhoods conditioned on observable characteristics and estimate conditional causal effects.
B.2 Simulation Evidence
In this section, we test the performance of the proposed method using the simulated dataset examined in the main text, where $N=400$ and $T=2$. We focus on the setting in which the units are embedded in a social network (top-left plot of Figure 2), generated by a random geometric graph (RGG) model (Penrose, 2003; Leung, 2022). Let $G_{ij}$ indicate the presence of an edge between units $i$ and $j$. Then $G_{ij}=\mathbf{1}\Big{\{}||\alpha_{i}-\alpha_{j}||\leq\sqrt{\frac{8}{\pi N}}\Big{\}}$, where $\alpha_{i}=(\alpha_{1i},\alpha_{2i})^{\prime}$ and each component is independently drawn from a uniform distribution on $[0,1]$. This specification yields a network $\mathcal{G}_{N}$ with an expected degree of 8.
We assume that there is one time-varying confounder in each $t\in\{1,2\}$ that takes the form of the outcome’s value in the previous period: $L_{i}^{(t)}=Y_{i}^{(t-1)}$, and the final outcome $Y_{i}=Y_{i}^{(2)}$. In the absence of any treatment, the potential values of $Y_{i}^{(t)}$ are generated from the following process:
$Y_{i}^{(t)}\left(\mathbf{0}_{\mathcal{N}}^{(1:2)}\right)=0.3+\xi^{(t)}+\varepsilon_{i}^{(t)},$
where $\xi^{(t)}$ is the period-specific intercept and $\varepsilon_{i}^{(t)}$ is the idiosyncratic error term for unit $i$ in period $t$. Both $\xi^{(t)}$ and $\varepsilon_{i}^{(t)}$ are independently drawn from the standard normal distribution. In each period $t$, unit $j$’s treatment generates a contemporaneous effect on unit $i$’s outcome with the magnitude of
$g\left(d_{ij},A_{i}^{(t)},A_{j}^{(t)}\right)=A_{j}^{(t)}\rho_{1i}\exp\left(-\frac{d_{ij}^{2}}{3}\right)+A_{i}^{(t)}A_{j}^{(t)}\rho_{2i}\exp\left(-\frac{d_{ij}^{2}}{3}\right),$
where $\rho_{1i}=\frac{\alpha_{1i}}{2}+1$ and $\rho_{2i}=\frac{\alpha_{2i}}{2}$ are unit-specific constants governing treatment effect heterogeneity. The effect’s magnitude depends on the proximity between two units and their treatment statues. It decreases exponentially as $d$ increases. We can verify that conditions required for Theorems 1 and 2 are satisfied in the data-generating process.
The effect received by each unit $i$ in period $t$ consists of four components: the average of contemporaneous effects from all the units in the sample, carryover effects from the previous period, influence of its own outcome (time-varying confounder) in the previous period, and the direct impacts of its first-degree neighbors’ outcome in the previous period (contagion):
$\tau_{i}^{(t)}\left(\mathbf{A}_{\mathcal{N}}^{(1:2)}\right)=$ $\underbrace{\sum_{d=0}^{d_{\max}}\frac{\sum_{j=1}^{N}\mathbf{1}\{d_{ij}=d\}g\left(d_{ij},A_{i}^{(t)},A_{j}^{(t)}\right)}{\sum_{j=1}^{N}\mathbf{1}\{d_{ij}=d\}}}_{\text{Contemporaneous effects}}+\underbrace{0.5*\tau_{i}^{(t-1)}\left(\mathbf{A}_{\mathcal{N}}^{(1:2)}\right)}_{\text{Carryover effects}}$
$+0.4*Y_{i}^{(t-1)}+\underbrace{0.1*\frac{\sum_{j=1}^{N}\mathbf{1}\{d_{ij}=1\}Y_{j}^{(t-1)}}{\sum_{j=1}^{N}\mathbf{1}\{d_{ij}=1\}}}_{\text{Congation}}.$
As mentioned in the main text, the carryover effect equals $50\%$ of the average contemporaneous effect in the previous period. Then, the observed value of unit $i$’s outcome in period $t$ equals
$Y_{i}^{(t)}=Y_{i}^{(t)}\left(\mathbf{A}_{\mathcal{N}}^{(1:2)}\right)=Y_{i}^{(t)}\left(\mathbf{0}_{\mathcal{N}}^{(1:2)}\right)+\tau_{i}^{(t)}\left(\mathbf{A}_{\mathcal{N}}^{(1:2)}\right).$
Structural equations in Section A.1 can be expressed as $\mathbf{L}_{i}^{(t)}=f_{\mathbf{L}}\left(\mathbf{A}_{\mathcal{N}_{\mathcal{G}_{N}}(i)}^{1:(t-1)},\mathbf{L}_{\mathcal{N}_{\mathcal{G}_{N}}(i)}^{1:(t-1)},\varepsilon_{\mathbf{L}i}^{(t)}\right)=Y_{i}^{(t-1)}\left(\mathbf{0}_{\mathcal{N}}^{(1:2)}\right)+\tau_{i}^{(t-1)}\left(\mathbf{A}_{\mathcal{N}}^{(1:2)}\right)$ and $Y_{i}=f_{Y}\left(\mathbf{A}_{\mathcal{N}_{\mathcal{G}_{N}}(i)}^{(1:T)},\mathbf{L}_{\mathcal{N}_{\mathcal{G}_{N}}(i)},\varepsilon_{Yi}\right)=Y_{i}^{(2)}\left(\mathbf{0}_{\mathcal{N}}^{(1:2)}\right)+\tau_{i}^{(2)}\left(\mathbf{A}_{\mathcal{N}}^{(1:2)}\right)$. ${\mathcal{N}_{\mathcal{G}_{N}}(i)}$’s form depends on the effect’s magnitude for each unit $i$, which is in turn decided by the idiosyncratic $\rho_{i}$.
In each period $t\in\{1,2\}$, unit $i$’s propensity score is generated by a logistical regression function of its own outcome and treatment status in the previous period:
$P(A_{i}^{(t)}=1)=Logit\left(-0.1+0.2*Y_{i}^{(t-1)}-0.6*A_{i}^{(t-1)}+\nu_{i}^{(t)}\right),$
where $\nu_{it}$ is drawn from a normal distribution with mean zero and standard deviation of $0.5$. Sequential exchangeability is clearly satisfied in this setting. As mentioned in the main text, we repeat the assignment process $1,000$ times and approximate each AMR using the average over the simulations. On average, the number of units under treatment histories $(0,0)$, $(1,0)$, $(0,1)$ and $(1,1)$ is $61$, $91$, $114$, and $134$, respectively.
For each unit $i$, its transformed outcome at each $d\in\{0,1,\dots,7\}$ is constructed as the average outcome across its $d$th-degree neighbors in the network. The propensity scores are estimated via logistic regressions fitted for each period. In each of the $1,000$ assignments, we rely on WLS to obtain $\hat{\boldsymbol{\beta}}(d)$ for each $d$ and plot them against their simulated true values in Figure B.2. Despite the moderate sample size ($400$ units), the bias is negligible, confirming the consistency of the proposed estimator.
Figure B.2: Bias of the Proposed Method
Notes: These plots show the bias of the WLS estimator for three combinations of MSM parameters. Black dots indicate the simulated true values of the causal effects at each proximity level, while gray dots represent estimates from each of the $1,000$ treatment assignments.
We further investigate how the performance of the method changes with an increasing sample size: $N\in\{400,625,900,1225,1600\}$. The top row of Figure B.3 shows the average mean squared error (MSE) of the estimates for each $d$ and $N$, while the bottom row displays the coverage rates. As the plots indicate, MSEs decline across all proximity levels as $N$ grows, and the coverage rates approach or exceed the nominal 95% level when the sample size is sufficiently large.
Figure B.3: Asymptotic Performance of the Proposed Method
Notes: The top panels display the average mean squared error (MSE) of the WLS estimator across $1,000$ treatment assignments, while the bottom panels show the coverage rates of the proposed 95% confidence intervals. Different colors correspond to different numbers of units used in estimation.
B.3 Extra Results from the Applications
To assess the validity of our identification assumptions in the two applications, we conduct placebo tests as described in Section A.5, estimating the effect of treatment history $(0,1)$ on outcomes in period 00. The main analysis assumes sequential exchangeability conditional on each unit’s treatment and outcomes from the previous period, implying that outcomes in period 00 should not be influenced by the treatment status in period $2$. The results are presented in Figure B.4. Consistent with our expectation, there is no significant effect on any of the four outcomes studied in the main text, thereby supporting the identification assumptions.
Figure B.4: Placebo Tests
(a) Placebo Tests in the Replication of Stokes (2016)
(b) Placebo Tests in the FHS Analysis
Notes: Panel (a) presents results from conducting placebo tests based on Stokes (2016), while Panel (b) displays the results from the FHS. Black and gray dots represent point estimates for different treatment histories and proximity levels. The vertical segments show the corresponding 95% confidence intervals, computed using the spatial/network HAC variance estimator and standard normal critical values.
Next, we examine the robustness of our findings to potential diffusion in treatment by modifying the propensity score model to incorporate the treatment statuses and time-varying confounders of nearby neighbors. In the replication of Stokes (2016), we include constituencies within a 5 km radius. For the FHS analysis, we consider up to three degrees of neighbors. Variables from each neighbor $j$ are weighted by the inverse of $d_{ij}$ when estimating the propensity score for unit $i$. Results are presented in Figure B.5. Most point estimates and their confidence intervals remain nearly identical to those obtained without adjusting for treatment diffusion.
Figure B.5: Results with Diffusion in Treatment
(a) Political Consequences of Proposed Wind Turbines
(b) Impacts of Smoking Cessation on Health Outcomes
Notes: Panel (a) presents results accounting for diffusion in treatment based on Stokes (2016), while Panel (b) displays the results from the FHS. Black and gray dots represent point estimates for different treatment histories and proximity levels. The vertical segments show the corresponding 95% confidence intervals, computed using the spatial/network HAC variance estimator and standard normal critical values.
Finally, we report effect estimates for three additional outcomes in the FHS: blood pressure, cholesterol level, and height. Among them, height serves as a placebo outcome, as it is unlikely to be influenced by smoking cessation. The results, presented in Figure B.6, suggest that smoking cessation leads to a moderate reduction in an individual’s cholesterol level but has little effect on their social network neighbors. In contrast, neither blood pressure nor height appears to be significantly affected by the treatment.
Figure B.6: mpacts of Smoking Cessation on Additional Health Outcomes
Notes: These plots present effect estimates for additional outcomes in the FHS. Black and gray dots represent point estimates for different treatment histories and proximity levels. The vertical segments show the corresponding 95% confidence intervals, computed using the network HAC variance estimator and standard normal critical values.